### v1

# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

Abstract
Reasoning is the fundamental capability of large language models (LLMs). Due to the rapid progress of LLMs, there are two main issues of current benchmarks: i) these benchmarks can be crushed in a short time (less than 1 year), and ii) these benchmarks may be easily hacked. To handle these issues, we propose the ever-scalingness for building the benchmarks which are uncrushable, unhackable, auto-verifiable and general. This paper presents Nondeterministic Polynomial-time Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. Specifically, the NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NP-complete problems and can generate any number of instances with any levels of complexities, ii) npsolver: which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, and iii) npeval: which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the aha moments, the reasoning errors and the solution errors. Extensive experiments over widely-used LLMs demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs’ performances to below 10%, demonstrating that NPPC is uncrushable, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1/o3-mini in most NP-complete problems considered, and iii) the numbers of tokens, aha moments in the advanced LLMs, e.g., Claude-3.7-Sonnet and DeepSeek-R1, are observed first to increase and then decrease when the problem instances become more and more difficult. We believe that NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for LLMs toward artificial general intelligence (AGI).
1 Introduction
Figure 1: Crush of Benchmarks
The remarkable successes of Large Language Models (LLMs) [1] have catalyzed the fundamental shift of artificial intelligence. The recent breakthrough on reasoning [14] enables the LLMs to complete complex tasks, e.g., math proof, code generation and computer use, which require the capabilities of understanding, generation and long-term planning. Various benchmarks, e.g., GPQA [24], AIME, SWE-bench [17] and ARC-AGI [6], are proposed to evaluate these advanced reasoning capabilities, where most benchmarks are curated and verified by human researchers with a finite number of questions. However, there are two main issues of current benchmarks: i) current benchmarks can be crushed in a short time (as shown in Figure 1): the performance on GSM8K [8] is increased from about 35% to 95% in three years, while the performance on SWE-bench [17] is increased from 7.0% to 64.6% in about 8 months, and ii) current benchmarks can be easily hacked or exploited: ChatbotArena leverages human votes to evaluate the LLMs, which may cost about \$ 3000 for one evaluation, while the MixEval [21] can reduce the cost to less than \$ 1 with high correlation. This low-cost evaluation brings the possibility of hacking ChatbotArena. The two main issues bring difficulties for the evaluation of LLMs and emerge as the main obstacles in a rapidly progressing era of LLMs.
Figure 2: Ever-scalingness
To build a successful benchmark for LLMs, we propose the ever-scalingness with the following four desiderata for a benchmark: i) uncrushable, which requires the scaling over the complexity, i.e., the benchmark can generate the problems with continually increasing complexities, ii) unhackable, which requires the scaling over instances, i.e., the benchmark can generate an infinite number of instances to avoid the exploitation, iii) auto-verifiable, which requires the scaling over the scalable oversight, i.e., the benchmark can verify the correctness of the solutions efficiently for the problems with any complexity, and iv) general, which requires the scaling over the coverage, i.e., the problems covered by the benchmark should be highly relevant to the real-world problems, rather than some puzzles or rare problems.
Figure 3: Motivation of NPPC
To build the ever-scaling benchmark, we consider nondeterminstic polynomial-time (NP) problems, i.e., the problems where the solutions can be verified in polynomial time [10]. Specifically, we focus on the NP-complete (NPC) problems, which are the hardest ones among all NP problems. The main reasons for choosing NPC problems for our ever-scaling benchmark are: i) the instances of NPC problems can be systematically generated with any level of difficulty, e.g., combinatorial search space, thus leading to the scaling of complexity and instances, ii) NPC problems are intrinsic to be “difficult to solve, easy to verify” and we do not have any polynomial-time algorithms to solve NPC problems. Therefore, the NPC problems are uncrushable even by any algorithms or tools so far and the solutions of any instance of NPC problems can be verified efficiently, i.e., scaling of the scalable oversight, and iii) many real-world problems, e.g., navigation and planning, can be formulated as NP(C) problems and any other NP problems can be transformed (or, reduced) to NPC problems in polynomial time. NPC problems are the foundation problems of all computational problems and LLMs are the foundation models for wide range tasks, thus leading to the emergence of our ever-scaling nondeterministic polynomial-time problem challenge (NPPC) (as displayed in Figure 3).
Specifically, the NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances with any levels of complexities, which implies the ever-scalingness of NPPC, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate users to evaluate their own models and iii) npeval, which provides comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the “aha moments”, the reasoning errors and the solution errors, which can provide in-depth analysis of the LLMs and the insights to further improve the LLMs’ reasoning capabilities. Extensive experiments over widely-used LLMs, i.e., GPT-4o-mini, GPT-4o, Claude-3.7-Sonnet, DeepSeek-V3, DeepSeek-R1, and OpenAI o1-mini, demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs’ performances to below 10%, demonstrating that NPPC is uncrushable, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1-mini in most NP-complete problems considered, and iii) the numbers of tokens, aha moments in the advanced LLMs, e.g.. Claude-3.7-Sonnet and DeepSeek-R1, are observed to first increase and then decrease when the problem instances become more and more difficult. We also analyze the different errors of reasoning and the solutions in the LLMs and the issues of offline models for solving complex reasoning problems. To the best of our knowledge, NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for the advanced LLMs toward artificial general intelligence (AGI). The code is released at https://github.com/SMU-DIGA/nppc.
2 Related Work
Table 1: Comparison of different reasoning benchmarks according to the ever-scalingness.
Uncrush.
Unhack.
Auto- verify
General
NPHardEval [13]
ZebraLogic [20]
Reasoning Gym [22]
Sudoku-Bench [25]
ARC-AGI-1 & 2 [6]
NPPC (this work)
We will provide a review of existing reasoning benchmarks in this section. Abstraction and Reasoning Corpus (ARC-AGI)-1 [6] is designed to be “easy for humans, hard for AI”, which is formed by human-curated 800 puzzle-like tasks, designed as grid-based visual reasoning problems. ARC-AGI-1 are featured by OpenAI as the leading benchmark to measure the performance of their o3 models. o3 at low compute scored 75.7% on ARC-AGI-1 and reached 87% accuracy with higher compute, which roughly crushes the ARC-AGI-1 benchmarks and leads to the emergence of the ARC-AGI-2 benchmark. ARC-AGI-1 & 2 are the representative of traditional benchmarks for LLMs, e.g., MMLU-Pro [28], GPQA [24], GSM8K [8], and SWE-bench [17], which are formed by static or regularly updated questions curated and verified by human. We believe that these benchmarks will never be uncrushable and may be hacked by some techniques.
Several recent benchmarks consider either NP(C) problems, e.g., 3SAT [4, 15, 23], or partially the ever-scalingness [13, 22], (displayed in Table 1). NPHardEval [13] considers 3 problems from P, NPC and NP-hard classes and use these class to evaluate the LLMs. We note that the problems in P class can be solved by augmenting the LLMs with tools, e.g., code running, and the NP-hard problems cannot be verified efficiently, therefore, NPHardEval cannot scale over the scalable oversight. Only 3 NPC problems are considered, i.e., Knapsack, Traveling salesman problem (TSP) and graph coloring, and the instances of each problem in NPHardEval are finite and only regularly updated, which cannot scale over the instance and complexity. ZebraLogic [20] considers one logic puzzle, i.e., Zebra puzzle, to test the reasoning capabilities of LMs when the problems’ complexities increase. However, the reasoning capability on specific puzzles does not necessarily transfer to other problems, which violates the scaling of the coverage. Sudoku-Bench [25] focuses on one specific Sudoku game with 2765 procedurally generated instances with various difficulty levels. Reasoning Gym [22] is an ongoing project which collects the procedural generators and algorithmic verifiers for infinite training data with adjustable complexity. Though with some NP(C) problems, e.g., Zebra puzzles and Sudoku, the reasoning gym does not specifically focus on NPC problems and scaling over complexity.
3 Preliminaries
3.1 P, NP and NP-complete Problems
Figure 4: Complexity classes
P and NP Problems.
The problems in P class are decision problems that can be solved in polynomial time by a deterministic Turing machine, which implies there exists an algorithm that can find a solution in time proportional to a polynomial function, e.g., $O(n^{k})$, of the input size $n$. Examples include sorting, shortest path problems, and determining if a number is prime. The problems in NP class are decision problems that can be solved in polynomial time by nondeterministic Turing machine, where a proposed solution can be easily verified, though finding that solution might require more time (as displayed in Definition 1). All P problems are also in NP, but the reverse remains an open question, known as “P vs. NP problem”. NP problems form the cornerstone of computational complexity theory, for which solution verification is tractable (polynomial time) even though solution discovery may be intractable (potentially exponential time), i.e., “difficult to solve, easy to verify”. Many real-world optimization problems can be formulated as NP problems, such as equilibrium finding in game theory, portfolio management, network design and machine learning.
Definition 1 (NP Problems).
The complexity class NP consists of all decision problems $\Omega$ such that for any “yes” instance $I$ of $\Omega$, there exists a certificate $\sigma$ of polynomial length in $|I|$ where a deterministic Turing machine can verify in polynomial time that $c$ is a valid certificate for $I$.
NP-complete (NPC) Problems.
Formally, a problem $\Omega$ is an NPC problem if i) the problem is in NP, and ii) any NP problems can be transformed to problem $\Omega$ in polynomial time. This reducibility property establishes NPC problems as the "hardest" problems in NP class. The Cook-Levin theorem established SAT as the first proven NPC problem [9, 18], while 3SAT is the special case of SAT and is also an NPC problem. Subsequent NPC problems typically proven via reduction chains back to 3SAT or other established NPC problems. The most well-known NPC problems include vertex cover problem, clique problem, traveling salesman proble (TSP), Hamiltonian path/cycle problem, etc. NPC problems play the most important roles in answering the “P vs. NP problem”, i.e., if any NPC problem were shown to have a polynomial-time algorithm, then P = NP. However, despite decades of research, no polynomial-time algorithms for any NPC problems have been discovered, which implies that the NPC problems are uncrushable by current methods or algorithms.
3.2 Reasoning in LLMs
The reasoning ability of LLMs refers to the model’s capacity to process information in a systematic way, which enables LLMs to tackle complex problems, e.g., mathematical proof and code generation, that require multi-step thinking, context understanding, and knowledge integration. Recently, specialized reasoning models have been proposed. OpenAI-o1 is an LLM trained with reinforcement learning (RL), which enables the model to perform complex reasoning, including logical thinking and problem solving, via chain-of-thought (CoT). o1 thinks before it answers and can significantly outperform GPT-4o on reasoning-heavy tasks with high data efficiency. DeepSeek-R1 [14] is an enhanced reasoning model designed to improve LLMs’ reasoning performance that incorporates multi-stage training and cold-start data before the large-scale RL. DeepSeek-R1 demonstrates remarkable reasoning capabilities, and achieves comparable performance to OpenAI-o1 across various reasoning tasks, such as mathematical problems, code generation, and scientific reasoning. Additionally, there are also open-sourced medium-sized LLMs with strong reasoning capabilities, e.g., DeepSeek-R1-32B, a distilled version of DeepSeek-R1, QwQ-32B [27], Gemma 3 [26].
4 Nondeterministic Polynomial-time Problem Challenge
Figure 5: Modules in NPPC
We introduce the Nondeterministic Polynomial Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. There are three main components in NPPC (as displayed in Figure 5): i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances and verify the solution with any levels of complexities, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate the users to evaluate their own models and iii) npeval, which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over problems, the number of tokens, the “aha moments”, the reasoning and solution errors, providing the in-depth analysis of the LLMs’ reasoning capabilities.
4.1 Problem Suite: npgym
Figure 6: Interaction loop between the LLM and the nygym.
Interaction Protocol.
[…]
Core Problems and Extension.
There are 25 typical NPC problems implemented in npgym. Among all NPC problems, we select a representative subset of 12 NPC problems as the core problems, ranging from the most famous NPC problems, e.g., 3SAT, Vertex Cover, and Clique, to the mathematical programming and string processing. The other 13 problems are categorized as the extension problems. A full list of the 25 problems is displayed in Table 2.
Table 2: Core Problems and Extension.
Problems
Core 3-Satisfiability (3SAT), Vertex Cover, 3-Dimensional Matching (3DM), Travelling Salesman (TSP), Hamiltonian Cycle, Graph 3-Colourability (3-COL), Bin Packing, Maximum Leaf Spanning Tree, Quadratic Diophantine Equations (QDE), Minimum Sum of Squares, Shortest Common Superstring, Bandwidth
Extension Clique, Independent Set, Dominating Set, Set Splitting, Set Packing, Exact Cover by 3-Sets (X3C), Minimum Cover, Partition, Subset Sum, Hitting String, Quadratic Congruences, Betweenness, Clustering
Generation and Verification.
Specifically, for each problem, npgym implements two functions:
  • generate_instance($\cdot$): given the configurations, this function will generate the problem instances. Taking the 3SAT as an example, the configurations include the number of variables and the number of clauses. The generated instances are guaranteed to have at least one solution and not necessarily to have a unique solution, which is ensured by the generation process.
  • verify_solution($\cdot$): given the solution and the problem instance, this function will verify whether the solution is correct or not. Additional to the correctness, this function also returns the error reasons. Taking the TSP as an example, the errors include i) the solution is not a tour, ii) the tour length exceeds the target length. The full list of the errors is displayed in Table 5.
Difficulty Levels.
Different NPC problems exhibit distinct combinatorial structures and computational characteristics. To establish a standardized metric for quantifying the computational complexity of these problems, npgym implements the difficulty levels [7, 13]. Each difficulty level corresponds to a specific parameterization. The current implementation of npgym stratifies each NPC problem into approximately 10 discrete difficulty levels, calibrated to induce a monotonically decreasing performance curve in LLMs, ranging from >90% success rate at level 1 to <10% at level 10. For each difficulty level, infinite number of problem instances can be generated, thus npgym is unhackable. The difficulty levels are empirically calibrated against current LLM performances and higher difficulty levels can be seamlessly included when the LLMs evolve to be more capable, which ensures that npgym is uncrushable. The full list of the difficulty levels is displayed in Appendix C.1.
4.2 Solver Suite: npsolver
Prompt Template.
The prompt template for LLMs is designed to be simple without any problem-specific knowledge and consistent across all problems. Therefore, the prompt template includes: i) problem description, which provides the concise definition of the NPC problem, including the problem name, the input and the question to be solved, ii) the context examples, where each example is formed by the instance and its corresponding solution, demonstrating the input and output patterns to help LLMs to generate the solution, iii) the target instance to solve, and iv) the general instruction about the solution format, where the solution is required to be in the JSON format for easy extracting and analyzing. We note that the structural output in JSON format may bring difficulties for LLMs to generate the correct solution, especially for the offline models, which will be analyzed in the experiments. The complete prompt template is displayed in Appendix D.
Completion with LLMs.
To streamline response extraction across various LLMs, we present npsolver, a solver suite that provides a unified interface for both online (API-based) and offline (locally deployed) models. npsolver includes: i) prompt generation, which constructs problem-specific prompts dynamically using the designed prompt templates, ii) LLM completion, that handles response generation via either online APIs supported through LiteLLM [5], or offline models via vLLM [19]; iii) solution extraction, which applies regular expressions to parse JSON-formatted responses, ensuring a consistent validation pipeline across all models; iv) error reporting, that standardizes error messages. Through the unified interface, npsolver enables both online and offline models to share a common workflow for completion and further evaluation.
4.3 Evaluation Suite: npeval
A comprehensive evaluation of the LLMs over all problems and all difficulty levels is time-consuming and expensive, mainly due to several sources of the randomness: i) the randomness of the generated instances, and ii) the randomness of the LLMs’ responses [^1] Randomizing the responses, i.e., non-zero temperature, is recommended for better performance [14].. Existing benchmarks let the LLMs solve the problem instances in a dataset, e.g., 200 instances in X-Large dataset in  [20], spanning across 5 difficulty levels and each level with 40 instances. However, we want to evaluate the LLMs’ performance on each difficulty level. Inspired by rliable [2], npeval considers the aggregated performance over different independent seeds, e.g., 3, for each difficulty level, and 30 instances are generated and evaluated for each seed, where 30 samples are generally considered as the minimum number for statistical analysis [16]. This sampling strategy enables statistically sound performance aggregation while controlling for instance-specific variance under the limited budget.
Figure 7: npeval.
Inspired by rliable [2], npeval provides the four measures to evaluate the performance: inter-quantile mean (IQM), mean, median, and the optimality gap and leverages stratified bootstrap confidence intervals (SBCIs) [11, 12] with stratified sampling for aggregate performance to report the interval estimation of the performance, which is a method that can be applied to small sample sizes and is better justified than reporting sample standard deviations. npeval provides the analysis of both prompt and completion tokens of LLMs across the problems and difficulty levels. As observed in [14], there are “aha moments” in the reasoning content. Therefore, npeval also provides the analysis of the number of “aha moments” during the reasoning. npeval provides the analysis of the errors including the errors of the generated solutions, i.e., the errors returned by npgym, and the reasoning errors, i.e., the errors in the internal reasoning process of LLMs for generating the solutions. This enables identification of the failures of LLMs. The main evaluations in npgym are displayed in Figure 7.
5 Results
Figure 8: Performance over difficulty levels
Figure 9: Ranks of models over problems
Figure 10: Performance interval over all problems across all levels
5.1 Analysis of Performance
The performance of online LLMs over difficulty levels is displayed in Figure 8, where all online models exhibit a decline in accuracy as difficulty levels increase across all 12 NPC problems. Take 3SAT as an example, all online models except for DeepSeek-R1 drop from $\geq 80\%$ accuracy to close to 0% at the last level, and DeepSeek-R1 shows the slowest decline but still falls to $\leq 15\%$ accuracy. All models collapse to around or even below $10\%$ accuracy at extreme difficulty confirms that NPPC is uncrushable against the SoTA LLMs and can discriminate their capabilities. One exception is Claude-3.7-Sonnet on Superstring problem, where the accuracy is still above 50% even for the level 10, while other models are all decreased into less than 20%, which demonstrates the superiority of Claude-3.7-Sonnet to deal with long contexts, where the prompts at level 10 is more than 50K [^2] We do not continually increase the difficulty of this problem as all other models are worse than 10%.. All models perform similarly on the Bandwidth problem, which may be mainly due to the fact that none of the models are familiar with this specific problem. Both o3-mini and DeepSeek-V3-2503 demonstrate superior performance to their predecessor models, o1-mini and DeepSeek-V3, respectively, validating continually improvements in both non-reasoning and reasoning LLMs.
The ranks of models over problems are shown in Figure 9, which measures the models’ performances across different levels of a specific problem. We observe that DeepSeek-R1 and o3-mini demonstrate statistical dominance in achievement of first-rank positions among reasoning-specialized architectures and Claude-3.7-Sonnet is the best non-reasoning model compared with the two versions of DeepSeek-v3 and GPT-4o, even better than o1-mini. Figure 10 visualizes the performance interval of different LLMs over all problems across all difficulty levels, where all four aggregate metrics are employed to measure LLMs’ performance. We observe that DeepSeek-R1 achieves superior performance with the highest IQM, mean, medium values and the lowest optimality gap, followed by o3-mini and Claude-3.7-Sonnet, while GPT-4o-mini performs in an opposite way.
5.2 Analysis of Tokens and Aha Moments
Tokens.
Figure 11 displays the token utilization across models on 3SAT. Offline models (QwQ-32B, DeepSeek-R1-32B) rapidly approach maximum token limits and incorrect solutions (red) usually take more tokens that correct solutions (blue). Among online models, DeepSeek-R1 demonstrates highest consumption (10,000-20,000 tokens) for successful solutions, while o-series models exhibit significant variance with outliers exceeding 40,000 tokens at higher complexity levels. DeepSeek-R1 and o3-mini show steeper token scaling compared to o1-mini and Claude-3.7-Sonnet, indicating advanced reasoning models leverage increased token allocation for complex problem-solving. GPT-4o variants maintain relatively efficient token utilization (<2,000) across all complexities. This quantifies the computational efficiency-performance tradeoff between specialized reasoning architectures and general-purpose models. Due to the limited space, full results over all problems are in Appendix I.
Figure 11: The number of tokens of different models on 3SAT
Aha Moments.
We also provide the analysis of the aha moments (instances of insight during reasoning, marked by phrases like “wait”) in the reasoning contents of DeepSeek-R1333The reasoning contents of o1/o3-mini are not available for analysis.. The results are displayed in Figure 12. While some problems (e.g., Superstring, TSP, and Bin Packing) show that the aha moments first increase and then decrease with difficulty, others exhibit steadier increases (QDE, 3DM) or relatively stable patterns (3SAT). This diversity suggests that advanced models’ approach to increasingly complex problems depends significantly on the specific problem domain and the necessities of investigating of different NPC problems.
Figure 12: Number of aha moments in DeepSeek-R1
5.3 Analysis of Solution Errors
Figure 13: The number of errors of different models on 3SAT
The solution errors of 3SAT is displayed in Figure  13. The results show that the distribution of these errors varies across models and difficulty levels. As the difficulty increases, the frequency of certain error types tends to increase as well. For QwQ-32B, JSON ERROR dominate across all levels, which is mainly due to the reasoning process is not finished when the context reaches the limits. For other models (such as GPT-4o, Claude-3.7-Sonnet, and DeepSeek-R1), problem-specific errors (3SAT ERROR 1 and 2) become more prevalent at higher difficulty levels. DeepSeek and GPT models show a more balanced distribution of error types. The o-series models demonstrate relatively better performance at intermediate difficulties but still encounter increasing problem-specific errors at the highest complexity levels. This error analysis provides valuable insights into not just whether models fail, but specifically how they fail when tackling increasingly complex NPC problems, which could inform future improvements to reasoning capabilities in LLMs.
5.4 Analysis of Reasoning Failure Cases
Table 3: Examples of Failure Cases for Deepseek-R1.
Failure Reason 1: Cascading Assumptions and Placeholder/Guess
… At this point, I’m not sure. Given the time I’ve spent and the complexity, perhaps the solution is the string pyeojmmykmadslpeetdb with length 20, even though some required substrings might be missing. But according to the problem statement, there is a solution. So I’ll have to take a guess. … But given the time I’ve spent and the need to provide an answer, I’ll make an educated guess. …
Failure Reason 2: Manual Computation Errors and Complexity Avoidance
So, if a node is in an edge that’s not covered by any other node, then that node must be in the vertex cover. But checking that for each edge would be time-consuming. Let’s look for edges that have a unique node. For example, if there’s an edge (a, b), and neither a nor b are connected to any other nodes, then one of them must be in the cover. But in a graph of 40 nodes, this is unlikely. So probably most nodes have multiple edges.
Failure Reason 3: Reliance on Prompt Examples
Wait, in the example problem, the solution included nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. So maybe similar patterns here. But the new problem’s edges are different. … Alternatively, perhaps the solution is similar to the example. The example’s solution includes nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. Looking at the problem’s edge set, perhaps some of these nodes are also high-degree here. Let’s see. … Wait, another idea: the example solution included nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. If I look at these nodes, perhaps they are all the nodes from one side of a bipartition. …
Failure Reason 4: Reasoning Repetition
… For example, node 8 appears in edges like (8,18), (8,35), (8,21), (8,30), (8,39), etc. Maybe node 8 is a hub. … In the problem to solve, node 8 is involved in many edges, as per the edges provided. So including 8 in the cover would cover a lot of edges. … For example, node 8 appears in a lot of edges. Including node 8 would cover many edges. … For example, many edges involve node 8. So including node 8 is a must.
For Deepseek-R1, the reasoning content of the failure cases shows several reasons that lead to wrong answers. i) cascading assumptions and placeholder/guess: DeepSeek-R1 begins with a high-level approach but quickly resorts to making assumptions to derive answers without logical deduction and considering all the conditions, and finally returns a placeholder or an educated guess; ii) manual computation errors and complexity avoidance: DeepSeek-R1 uses inefficient manual calculations (prone to errors) instead of programming, skips complex steps even the reasoning is correct, and resorts to guesses to avoid effort; iii) reliance on prompt examples: DeepSeek-R1 relies heavily on the example solution, making it waste time and get distracted by verifying and editing the solution instead of solving the problem directly; iv) reasoning repetition: DeepSeek-R1 gets stuck repeating the same logic without making further progress, wasting time and tokens. We list some typical examples of failure cases of DeepSeek-R1 in Table 3, and more examples are shown in Table 20 in Appendix L. For Claude-3.7-Sonnet, its failure cases typically exhibit more concise reasoning. Claude often outlines a high-level step-by-step approach but omits detailed calculations and rigorous verification, and it relies on approximate calculations to derive a final answer, incorrectly asserting that the result has been validated. Example is shown in Table 21 in Appendix L.
6 Conclusions
Reasoning stands as the foundational capability of large language models (LLMs). However, the rapid advancement of LLMs’ reasoning abilities has rendered current benchmarks easily crushable and vulnerable to hacking. Therefore, we propose Nondeterministic Polynomial Problem challenge (NPPC), an ever-scaling benchmark that is uncrushable, unhackable, auto-verifiable, and general, designed to evolve alongside LLM advancements. NPPC comprises three core components: i) npgym: a unified framework for generating customizable problem instances across 25 NPC problems with adjustable complexity levels; ii) npsolver: a flexible evaluation interface supporting both online APIs and offline local deployments; iii) npeval: a comprehensive toolkit for the systematic evaluation of LLMs across different problems, including the solution validity, reasoning errors, token efficiency. Our extensive experiments with state-of-the-art LLMs demonstrate that: i) NPPC successfully reduces all models’ performance to below 10% at extreme difficulties, confirming its uncrushable nature, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1-mini emerge as the most powerful LLMs, with DeepSeek-R1 outperforming others in 7/12 problems, iii) Models exhibit distinct failure patterns, including cascading assumptions, manual computation errors, and reasoning repetition. To the best of our knowledge, NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for LLMs toward artificial general intelligence (AGI).
Appendix
Appendix A Frequently Asked Questions (FAQs)
A.1 Why Ever-Scaling and Why the Four Properties are Important?
Why Ever-Scaling?
LLMs are advancing at an unprecedented pace, making existing benchmarks obsolete quickly and posing a significant challenge for maintaining reliable evaluation. An ever-scaling benchmark can evolve alongside LLMs, i.e., adapting dynamically to match the development of LLMs. The ever-scaling benchmark can address two core limitations in traditional benchmarks: i) short lifespan, where traditional benchmarks are easily crushed as LLMs rapidly improve, losing their ability to distinguish between models; ii) limited exploitability, where models can hack the answers in static benchmarks through overfitting or finding shortcuts to answers without genuine reasoning.
Why the Four Properties are Important?
The four properties include:
  • Uncrushable (scaling over the complexity): The benchmark can generate problems with continually increasing difficulty, e.g., larger input sizes, stricter constraints, etc. This property can prevent the benchmark from being solved to prevent obsolescence, and mirror the real-world problems, e.g., logistics and chip design, which grow in complexity as systems scale.
  • Unhackable (scaling over the instances): The benchmark can generate infinite unique instances, even at the same complexity level. This unhackable property makes it impossible for LLMs to memorize the answers or simply overfit to patterns in static training data, and it forces LLMs to reason about the underlying logic to ensure the fairness of evaluation.
  • Auto-verifiable (scaling over the scalable oversight): The benchmark provides an automated and cost-effective evaluation without any human intervention, i.e., the solutions can be verified efficiently in polynomial time even for arbitrarily complex problems. This property is critical for large-scale benchmarking as human evaluation is impractical for massive or highly complex benchmarks, therefore, automated verification is necessary for evaluating at scale.
  • General (scaling over the coverage): This property ensures the benchmark to focus on problems with broad applicability to reflect real-world utility and challenges, enabling the progress made on benchmarks to indicate the progress made on real tasks.
A.2 Why Focusing on NP (Specifically NPC) Problems?
Why not P or NP-hard problems?
The problems in P can be solved in polynomial time. If the LLM is equipped with the tools to run the code, it can generate the code to solve these problems. In this case, the benchmark can be easily crushed by the tool using. While for NP-hard problems, especially for the problems that cannot be verified in polynomial time, when the problems become extremely large, we cannot verify the solution efficiently, which may hurt the scaling over complexity.
Why NPC Problems?
NPC problems are the “hardest” problems in NP class and any other NP problems can be reduced to NPC problems in polynomial time. The absence of known polynomial-time algorithms for NPC problems ensures that current benchmarks measuring performance on these problems cannot be trivially dominated through tool using. Furthermore, the polynomial-time verifiability of solutions enables efficient assessment of solutions generated by LLMs or AI agents even for large problem instances.
A.3 Why Not Considering More Complex Test-time Scaling?
The Majority Voting, Best of $N$, and even tools, e.g., domain-specific solvers, can further improve the performance of models [23, 20]. However, these approaches either necessitate multiple forward passes through the language model or incorporate auxiliary components such as reward models or external tools to augment the reasoning process. Our primary objective is to investigate the reasoning capabilities of LLMs and these complex test-time scaling would be beyond the scope of this paper.
A.4 Why More NP-complete Problems are Needed, i.e., Why Not Focusing on 3SAT Only?
3SAT is a classic NPC problem with theoretical completeness, which provides a theoretically rigorous foundation for benchmarking. As an NPC problem, although all NP problems can be reduced to 3SAT, solely relying on reduction to 3SAT is impractical and reasoning benchmarks demand broader diversity for several key reasons:
  • Reduction overhead: The reduction process may incur significant computational overhead. Additional variables and constraints are often introduced when reducing non-trivial NP problems to a specific NP-complete problem, e.g., reducing Traveling Salesman Problem (TSP) to 3SAT requires mapping the structure of the original problem into a Boolean logic expression through an encoding mechanism, which introduces an exponential number of variables and clauses, significantly increasing the computational complexity.
  • Loss of characteristics: Each specific NP problem has domain-specific information, e.g., structure and characteristics. For example, Traveling Salesman Problem (TSP) has graph structures, Bin Packing has combinatorial optimization characteristics, and Graph 3-Colourability (3-COL) has adjacency characteristics. Therefore, reducing NP problems to 3SAT and only considering 3SAT will cause the loss of problem specificity, e.g., structural semantics, which could be used to design more efficient heuristics or approximation algorithms.
  • Lack of robustness: NP problems form the foundation of numerous real-world scenarios, which often exhibit various conditions that cannot be adequately represented solely through 3SAT. As a reasoning benchmark, NPPC should encompass a variety of problem sizes and structures rather than concentrating exclusively on 3SAT to effectively evaluate the capabilities and scalability of LLMs. Therefore, a diverse set of complex NP problems that can closely mimic real-world challenges should be considered.
A.5 Discussion about Future Works
Unstoppable RL vs. Ever-Scaling NP Problems. The rapid progress in LLM reasoning capabilities through reinforcement learning (RL) presents an interesting dynamic when considered alongside ever-scaling NPC problems. As models like DeepSeek-R1 and OpenAI o1/o3-mini demonstrate significant reasoning improvements through RL techniques, NPPC provides a counterbalance by offering problems that can continuously scale in difficulty. This creates an adversarial paradigm to drive the AI development: RL improves model reasoning and NPPC scales to maintain challenging.
Multimodal NP Problems. Extending NPPC to the multimodal domains represents a promising direction. Games like StarCraft II, Minesweeper, Pokemon and Super Mario Bros [3], could form the foundation of a multimodal version of NPPC. This would enable testing reasoning capabilities across multiple modalities while maintaining the ever-scaling properties. A multimodal NPPC could evaluate how well models can reason about visual, spatial, and temporal information jointly.
AI Agent. The benchmark could significantly contribute to AI agent development by encouraging tool use for solving increasingly complex NP problems. As the difficulty of problems increases, LLMs will naturally require external tools to manage computational complexity. This creates a natural pathway toward agent capabilities, where models learn to decompose problems and leverage appropriate tools. The code generation already observed in models attempting to solve difficult NPPC problems can be viewed as a form of tool creation, as these generated algorithms can be saved and reused for future problem-solving. This provides a principled way to measure progress in agent development within a well-defined formal framework.
Appendix B Computational Complexity: P, NP and NP-complete
Figure 14: The relation between P, NP and NP-complete
P.
The class P consists of decision problems that can be solved by a deterministic Turing machine in polynomial time. In practical terms, these are problems for which efficient algorithms exist. The time required to solve these problems grows polynomially with the input size ($n$), such as $O(n)$, $O(n^{2})$, or $O(n^{3})$. Examples include sorting, searching in a sorted array, and determining if a number is prime.
NP.
NP contains all decision problems for which a solution can be verified in polynomial time. Every problem in P is also in NP, but NP may contain problems that are not in P. The key characteristic is that if someone gives you a potential solution, you can quickly check whether it’s correct, even if finding that solution might be difficult. Examples include the Boolean satisfiability problem and the Traveling Salesman decision problem.
NP-complete.
NP-complete problems are the “hardest” problems in NP. A problem is NP-complete if: i) It belongs to NP, ii) Every other problem in NP can be reduced to it in polynomial time. This means that if an efficient (polynomial-time) algorithm were found for any NP-complete problem, it could be used to solve all problems in NP efficiently. The first proven NP-complete problem was the Boolean satisfiability problem (SAT). Other examples include the Traveling Salesman Problem, Graph Coloring, and the Knapsack Problem. The question of whether P=NP (whether every problem with efficiently verifiable solutions also has efficiently computable solutions) remains one of the most important open questions in computer science and mathematics.
Appendix C Modules in NPPC
C.1 Problem Suite: npgym
Interface.
We introduce npgym, a problem suite containing 25 NPC problems with a unified gym-style interface for instance generation and solution verification. Each environment is defined by a problem name and its corresponding hyperparameters, enabling the generation of unlimited problem instances and example solutions. Difficulty can be scaled by adjusting these parameters. npgym also supports automatic verification of solutions produced by large language models (LLMs). New problems can be added easily by implementing two core functions and providing a problem description for prompt generation.
class NPEnv:
def __init__(self, problem_name, level):
self.problem_name = problem_name
self.level = level
self._generate_instance, self._verify_solution = self._get_instance_generator()
def _get_instance_generator(self):
np_gym_folder = "./npgym/npc"
problem_path = PROBLEM2PATH[self.problem_name]
generate_instance = importlib.import_module(problem_path).generate_instance
verify_solution = importlib.import_module(problem_path).verify_solution
return generate_instance, verify_solution
Variables to Scale.
Table 4 lists the variables to scale for each of the 25 NP-complete problems.
Table 4: NPC problems in NPPC and the variables to scale
Type Problems Variables to scale
Core 3SAT num_variables, num_clauses
Vertex Cover num_nodes, cover_size
3DM n
TSP num_cities, target_length
Hamiltonian Cycle num_nodes, directed
3-COL num_nodes, num_edges
Bin Packing num_items, bin_capacity, num_bins
Max Leaf Span Tree num_nodes, target_leaves
QDE low, high
Min Sum of Squares num_elements, k
Superstring n, k
Bandwidth num_nodes, bandwidth
Extension Clique num_nodes, clique_size
Independent Set num_nodes, ind_set_size
Dominating Set num_nodes, k, edge_prob
Set Splitting num_elements, num_subsets
Set Packing num_elements, num_subsets, num_disjoint_sets
X3C num_elements, num_subsets
Minimum Cover num_elements, num_sets, k
Partition n, max_value
Subset Sum num_elements, max_value
Hitting String n, m
Quadratic Congruences min_value, max_value
Betweenness num_element, num_triples
Clustering num_elements, b
Difficulty Levels.
We define and release problem-specific difficulty levels for each of the 25 core problems included in our benchmark. Each problem includes approximately 10 levels of increasing complexity, determined primarily by theoretical factors such as search space size and validated through empirical testing using DeepSeek-R1 and GPT-4o. npgym allows seamless extension to higher difficulty levels as more powerful models become available.
"3-Satisfiability (3-SAT)": {
1: {"num_variables": 5, "num_clauses": 5},
2: {"num_variables": 15, "num_clauses": 15},
3: {"num_variables": 20, "num_clauses": 20},
4: {"num_variables": 25, "num_clauses": 25},
5: {"num_variables": 30, "num_clauses": 30},
6: {"num_variables": 40, "num_clauses": 40},
7: {"num_variables": 50, "num_clauses": 50},
8: {"num_variables": 60, "num_clauses": 60},
9: {"num_variables": 70, "num_clauses": 70},
10: {"num_variables": 80, "num_clauses": 80},
"Vertex Cover": {
1: {"num_nodes": 4, "cover_size": 2},
2: {"num_nodes": 8, "cover_size": 3},
3: {"num_nodes": 12, "cover_size": 4},
4: {"num_nodes": 16, "cover_size": 5},
5: {"num_nodes": 20, "cover_size": 10},
6: {"num_nodes": 24, "cover_size": 12},
7: {"num_nodes": 28, "cover_size": 14},
8: {"num_nodes": 32, "cover_size": 16},
9: {"num_nodes": 36, "cover_size": 18},
10: {"num_nodes": 40, "cover_size": 20},
"Clique": {
1: {"num_nodes": 4, "clique_size": 2},
2: {"num_nodes": 8, "clique_size": 4},
3: {"num_nodes": 12, "clique_size": 6},
4: {"num_nodes": 14, "clique_size": 7},
5: {"num_nodes": 16, "clique_size": 8},
6: {"num_nodes": 18, "clique_size": 9},
7: {"num_nodes": 20, "clique_size": 10},
8: {"num_nodes": 22, "clique_size": 11},
9: {"num_nodes": 24, "clique_size": 12},
10: {"num_nodes": 26, "clique_size": 13},
11: {"num_nodes": 28, "clique_size": 14},
12: {"num_nodes": 30, "clique_size": 15},
13: {"num_nodes": 40, "clique_size": 20},
"Independent Set": {
1: {"num_nodes": 4, "ind_set_size": 2},
2: {"num_nodes": 8, "ind_set_size": 4},
3: {"num_nodes": 12, "ind_set_size": 6},
4: {"num_nodes": 16, "ind_set_size": 8},
5: {"num_nodes": 20, "ind_set_size": 10},
6: {"num_nodes": 24, "ind_set_size": 12},
7: {"num_nodes": 26, "ind_set_size": 13},
8: {"num_nodes": 28, "ind_set_size": 14},
9: {"num_nodes": 30, "ind_set_size": 15},
10: {"num_nodes": 32, "ind_set_size": 16},
11: {"num_nodes": 34, "ind_set_size": 17},
12: {"num_nodes": 36, "ind_set_size": 18},
13: {"num_nodes": 48, "ind_set_size": 24},
"Partition": {
1: {"n": 2, "max_value": 1},
2: {"n": 4, "max_value": 40},
3: {"n": 10, "max_value": 100},
4: {"n": 20, "max_value": 200},
5: {"n": 30, "max_value": 300},
6: {"n": 40, "max_value": 400},
7: {"n": 50, "max_value": 500},
8: {"n": 55, "max_value": 550},
9: {"n": 60, "max_value": 600},
10: {"n": 65, "max_value": 650},
11: {"n": 70, "max_value": 700},
12: {"n": 75, "max_value": 750},
13: {"n": 80, "max_value": 800},
"Subset Sum": {
1: {"num_elements": 5, "max_value": 100},
2: {"num_elements": 10, "max_value": 100},
3: {"num_elements": 20, "max_value": 200},
4: {"num_elements": 40, "max_value": 400},
5: {"num_elements": 80, "max_value": 800},
6: {"num_elements": 100, "max_value": 1000},
7: {"num_elements": 120, "max_value": 1200},
8: {"num_elements": 160, "max_value": 1000},
9: {"num_elements": 160, "max_value": 1600},
10: {"num_elements": 200, "max_value": 2000},
11: {"num_elements": 200, "max_value": 1000},
12: {"num_elements": 400, "max_value": 2000},
13: {"num_elements": 600, "max_value": 2000},
"Set Packing": {
1: {"num_elements": 10, "num_subsets": 10, "num_disjoint_sets": 2},
2: {"num_elements": 40, "num_subsets": 40, "num_disjoint_sets": 8},
3: {"num_elements": 100, "num_subsets": 200, "num_disjoint_sets": 50},
4: {"num_elements": 100, "num_subsets": 400, "num_disjoint_sets": 30},
5: {"num_elements": 100, "num_subsets": 500, "num_disjoint_sets": 30},
6: {"num_elements": 100, "num_subsets": 600, "num_disjoint_sets": 30},
7: {"num_elements": 100, "num_subsets": 800, "num_disjoint_sets": 30},
8: {"num_elements": 100, "num_subsets": 1000, "num_disjoint_sets": 30},
9: {"num_elements": 200, "num_subsets": 400, "num_disjoint_sets": 60},
10: {"num_elements": 200, "num_subsets": 800, "num_disjoint_sets": 60},
11: {"num_elements": 400, "num_subsets": 1000, "num_disjoint_sets": 200},
"Set Splitting": {
1: {"num_elements": 5, "num_subsets": 5},
2: {"num_elements": 10, "num_subsets": 10},
3: {"num_elements": 10, "num_subsets": 50},
4: {"num_elements": 10, "num_subsets": 100},
5: {"num_elements": 10, "num_subsets": 200},
6: {"num_elements": 100, "num_subsets": 100},
7: {"num_elements": 100, "num_subsets": 200},
8: {"num_elements": 10, "num_subsets": 500},
9: {"num_elements": 10, "num_subsets": 1000},
10: {"num_elements": 15, "num_subsets": 500},
11: {"num_elements": 20, "num_subsets": 500},
"Shortest Common Superstring": {
1: {"n": 10, "k": 5},
2: {"n": 20, "k": 10},
3: {"n": 40, "k": 20},
4: {"n": 80, "k": 40},
5: {"n": 100, "k": 50},
6: {"n": 100, "k": 100},
7: {"n": 100, "k": 200},
8: {"n": 200, "k": 200},
9: {"n": 300, "k": 400},
10: {"n": 300, "k": 600},
"Quadratic Diophantine Equations": {
1: {"low": 1, "high": 50},
2: {"low": 1, "high": 100},
3: {"low": 1, "high": 500},
4: {"low": 1, "high": 1000},
5: {"low": 1, "high": 5000},
6: {"low": 1, "high": 10000},
7: {"low": 1, "high": 50000},
8: {"low": 1, "high": 80000},
9: {"low": 1, "high": 100000},
10: {"low": 1, "high": 200000},
"Quadratic Congruences": {
1: {"min_value": 1, "max_value": 100},
2: {"min_value": 1, "max_value": 1000},
3: {"min_value": 1, "max_value": 10000},
4: {"min_value": 1, "max_value": 50000},
5: {"min_value": 1, "max_value": 100000},
6: {"min_value": 1, "max_value": 300000},
7: {"min_value": 1, "max_value": 500000},
8: {"min_value": 1, "max_value": 800000},
9: {"min_value": 1, "max_value": 1000000},
10: {"min_value": 1, "max_value": 3000000},
"3-Dimensional Matching (3DM)": {
1: {"n": 4},
2: {"n": 8},
3: {"n": 12},
4: {"n": 15},
5: {"n": 20},
6: {"n": 25},
7: {"n": 30},
8: {"n": 40},
9: {"n": 50},
10: {"n": 60},
"Travelling Salesman (TSP)": {
1: {"num_cities": 5, "target_length": 100},
2: {"num_cities": 8, "target_length": 100},
3: {"num_cities": 10, "target_length": 100},
4: {"num_cities": 12, "target_length": 100},
5: {"num_cities": 15, "target_length": 100},
6: {"num_cities": 17, "target_length": 200},
7: {"num_cities": 20, "target_length": 200},
8: {"num_cities": 25, "target_length": 200},
9: {"num_cities": 30, "target_length": 200},
10: {"num_cities": 40, "target_length": 300},
"Dominating Set": {
1: {"num_nodes": 10, "k": 5, "edge_prob": 0.3},
2: {"num_nodes": 15, "k": 5, "edge_prob": 0.3},
3: {"num_nodes": 30, "k": 15, "edge_prob": 0.3},
4: {"num_nodes": 50, "k": 20, "edge_prob": 0.3},
5: {"num_nodes": 70, "k": 20, "edge_prob": 0.3},
6: {"num_nodes": 100, "k": 20, "edge_prob": 0.3},
7: {"num_nodes": 70, "k": 20, "edge_prob": 0.2},
8: {"num_nodes": 80, "k": 20, "edge_prob": 0.2},
9: {"num_nodes": 100, "k": 20, "edge_prob": 0.2},
10: {"num_nodes": 150, "k": 20, "edge_prob": 0.2},
11: {"num_nodes": 160, "k": 15, "edge_prob": 0.2},
12: {"num_nodes": 180, "k": 15, "edge_prob": 0.2},
"Hitting String": {
1: {"n": 5, "m": 10},
2: {"n": 5, "m": 20},
3: {"n": 10, "m": 20},
4: {"n": 10, "m": 30},
5: {"n": 10, "m": 40},
6: {"n": 10, "m": 45},
7: {"n": 10, "m": 50},
8: {"n": 10, "m": 55},
9: {"n": 10, "m": 60},
10: {"n": 10, "m": 70},
"Hamiltonian Cycle": {
1: {"num_nodes": 5, "directed": False},
2: {"num_nodes": 8, "directed": False},
3: {"num_nodes": 10, "directed": False},
4: {"num_nodes": 12, "directed": False},
5: {"num_nodes": 16, "directed": False},
6: {"num_nodes": 18, "directed": False},
7: {"num_nodes": 20, "directed": False},
8: {"num_nodes": 22, "directed": False},
9: {"num_nodes": 25, "directed": False},
10: {"num_nodes": 30, "directed": False},
"Bin Packing": {
1: {"num_items": 10, "bin_capacity": 20, "num_bins": 3},
2: {"num_items": 20, "bin_capacity": 30, "num_bins": 3},
3: {"num_items": 30, "bin_capacity": 30, "num_bins": 3},
4: {"num_items": 40, "bin_capacity": 30, "num_bins": 3},
5: {"num_items": 50, "bin_capacity": 50, "num_bins": 5},
6: {"num_items": 60, "bin_capacity": 50, "num_bins": 5},
7: {"num_items": 70, "bin_capacity": 50, "num_bins": 5},
8: {"num_items": 80, "bin_capacity": 50, "num_bins": 5},
9: {"num_items": 80, "bin_capacity": 30, "num_bins": 10},
10: {"num_items": 100, "bin_capacity": 50, "num_bins": 10},
"Exact Cover by 3-Sets (X3C)": {
1: {"num_elements": 3, "num_subsets": 6},
2: {"num_elements": 4, "num_subsets": 8},
3: {"num_elements": 5, "num_subsets": 10},
4: {"num_elements": 7, "num_subsets": 14},
5: {"num_elements": 8, "num_subsets": 16},
6: {"num_elements": 10, "num_subsets": 20},
7: {"num_elements": 15, "num_subsets": 30},
8: {"num_elements": 20, "num_subsets": 40},
9: {"num_elements": 25, "num_subsets": 50},
10: {"num_elements": 30, "num_subsets": 60},
"Minimum Cover": {
1: {"num_elements": 5, "num_sets": 10, "k": 3},
2: {"num_elements": 10, "num_sets": 20, "k": 5},
3: {"num_elements": 10, "num_sets": 30, "k": 5},
4: {"num_elements": 15, "num_sets": 20, "k": 8},
5: {"num_elements": 15, "num_sets": 30, "k": 10},
6: {"num_elements": 20, "num_sets": 40, "k": 10},
7: {"num_elements": 25, "num_sets": 50, "k": 10},
8: {"num_elements": 30, "num_sets": 60, "k": 10},
9: {"num_elements": 35, "num_sets": 70, "k": 10},
10: {"num_elements": 40, "num_sets": 80, "k": 10},
11: {"num_elements": 45, "num_sets": 90, "k": 10},
12: {"num_elements": 50, "num_sets": 100, "k": 10},
13: {"num_elements": 55, "num_sets": 110, "k": 10},
14: {"num_elements": 60, "num_sets": 120, "k": 10},
15: {"num_elements": 65, "num_sets": 130, "k": 10},
16: {"num_elements": 70, "num_sets": 140, "k": 10},
"Graph 3-Colourability (3-COL)": {
1: {"num_nodes": 5, "num_edges": 8},
2: {"num_nodes": 8, "num_edges": 12},
3: {"num_nodes": 10, "num_edges": 20},
4: {"num_nodes": 15, "num_edges": 25},
5: {"num_nodes": 15, "num_edges": 30},
6: {"num_nodes": 15, "num_edges": 40},
7: {"num_nodes": 20, "num_edges": 40},
8: {"num_nodes": 20, "num_edges": 45},
9: {"num_nodes": 30, "num_edges": 60},
10: {"num_nodes": 30, "num_edges": 80},
"Clustering": {
1: {"num_elements": 6, "b": 10},
2: {"num_elements": 10, "b": 10},
3: {"num_elements": 15, "b": 10},
4: {"num_elements": 18, "b": 10},
5: {"num_elements": 20, "b": 10},
6: {"num_elements": 30, "b": 10},
7: {"num_elements": 40, "b": 10},
8: {"num_elements": 50, "b": 10},
9: {"num_elements": 60, "b": 10},
10: {"num_elements": 70, "b": 10},
"Betweenness": {
1: {"num_element": 3, "num_triples": 1},
2: {"num_element": 4, "num_triples": 2},
3: {"num_element": 5, "num_triples": 3},
4: {"num_element": 6, "num_triples": 4},
5: {"num_element": 7, "num_triples": 5},
6: {"num_element": 8, "num_triples": 6},
"Minimum Sum of Squares": {
1: {"num_elements": 10, "k": 5},
2: {"num_elements": 50, "k": 8},
3: {"num_elements": 100, "k": 8},
4: {"num_elements": 100, "k": 5},
5: {"num_elements": 100, "k": 4},
6: {"num_elements": 100, "k": 3},
7: {"num_elements": 200, "k": 10},
8: {"num_elements": 200, "k": 4},
9: {"num_elements": 200, "k": 3},
10: {"num_elements": 300, "k": 3},
"Bandwidth": {
1: {"num_nodes": 3, "bandwidth": 2},
2: {"num_nodes": 4, "bandwidth": 2},
3: {"num_nodes": 5, "bandwidth": 3},
4: {"num_nodes": 6, "bandwidth": 3},
5: {"num_nodes": 5, "bandwidth": 2},
6: {"num_nodes": 7, "bandwidth": 3},
7: {"num_nodes": 6, "bandwidth": 2},
8: {"num_nodes": 8, "bandwidth": 3},
9: {"num_nodes": 7, "bandwidth": 2},
10: {"num_nodes": 8, "bandwidth": 2},
"Maximum Leaf Spanning Tree": {
1: {"num_nodes": 5, "target_leaves": 2},
2: {"num_nodes": 10, "target_leaves": 5},
3: {"num_nodes": 20, "target_leaves": 10},
4: {"num_nodes": 30, "target_leaves": 20},
5: {"num_nodes": 40, "target_leaves": 30},
6: {"num_nodes": 60, "target_leaves": 50},
7: {"num_nodes": 70, "target_leaves": 60},
8: {"num_nodes": 80, "target_leaves": 65},
9: {"num_nodes": 90, "target_leaves": 75},
10: {"num_nodes": 100, "target_leaves": 80},
Solution Errors.
There are two fundamental error categories: problem-independent errors and problem-dependent errors. Problem-independent errors are general errors that arise from external factors unrelated to the problem’s intrinsic characteristics and all problems have these types of errors. Problem-independent errors include JSON ERROR (JSON not found or JSON parsing errors), and VERIFICATION ERROR (output format mismatches or structural validation failures). Problem-dependent errors originate from the problem’s inherent complexity, which are defined based on problem specificity. A comprehensive illustration of the errors is displayed in Table 5.
Table 5: A comprehensive illustration of errors.
Problem Error Type Description
JSON ERROR JSON not found.
VERIFICATION ERROR Wrong output format.
3SAT ERROR 1 The solution length mismatches the number of variables.
ERROR 2 Some clauses are not satisfied.
Vertex Cover ERROR 1 Wrong solution format.
ERROR 2 The cover is empty.
ERROR 3 Invalid vertex index, i.e., above the max or below the min.
ERROR 4 The cover size exceeds the limit.
ERROR 5 Some edges are not covered.
3DM ERROR 1 Not all triples in the matching are in the original set.
ERROR 2 The size of matching is wrong
ERROR 3 The elements in the matching are not mutually exclusive.
TSP ERROR 1 Tour length mismatches number of cities.
ERROR 2 Invalid city index, i.e., above the max or below the min.
ERROR 3 There exists cities not be visited exactly once.
ERROR 4 Tour length exceeds target length.
Hamiltonian Cycle ERROR 1 Path length is wrong.
ERROR 2 Path does not return to start.
ERROR 3 Not all vertices visited exactly once.
ERROR 4 There exists invalid vertex in path.
ERROR 5 There exists invalid edges in path.
3-COL ERROR 1 The two nodes of an edge have the same color
Bin Packing ERROR 1 Solution length mismatches the number of items.
ERROR 2 Invalid bin index.
ERROR 3 The total size exceeds bin capacity.
Max Leaf Span Tree ERROR 1 Solution length mismatches the number of vertices.
ERROR 2 There exists invalid edges in solution.
ERROR 3 The solution does not have exactly one root.
ERROR 4 The solution doesn’t span all vertices.
ERROR 5 The number of leaves in the solution is less than target.
QDE ERROR 1 Solution length mismatches the number of integers.
ERROR 2 There exists non-positive values in the solution.
ERROR 3 The equation does not hold.
Min Sum Square ERROR 1 Solution length mismatches the number of elements.
ERROR 2 The number of subsets exceeds the set limit.
ERROR 3 The sum exceeds the limit $J$.
Superstring ERROR 1 Wrong solution format.
ERROR 2 The solution length exceeds the limit.
ERROR 3 Some string is not the substring of the solution.
Bandwidth ERROR 1 Layout length mismatches the number of vertices.
ERROR 2 Layout is not a permutation of vertices.
ERROR 3 There exists edge exceeds the bandwidth limit.
C.2 Solver Suite: npsolver
We introduce npsolver, a solver suite that provides a unified interface for both online (API-based) and offline (local) models. The unified interface includes: i) Prompt Generation, which constructs problem-specific prompts dynamically using the designed prompt templates shown in Appendix D, including problem descriptions, in-context examples, and target problems; ii) LLM Completion, which invokes either online or offline LLMs to generate responses from the constructed prompts; iii) Solution Extraction, which designs regular expressions to parse JSON outputs from the LLMs’ responses, ensuring all online and offline LLMs Use the same JSON validation pipeline; iv) Error Reporting, which standardizes error messages. Through the unified interface, npsolver enables both online and offline models to share a common workflow. Through this unified pipeline, npsolver enables consistent evaluation and analysis for both online and offline models. For each problem, difficulty level, and model, npsolver stores detailed records—including the problem instance, example solutions, full LLM responses, extracted solutions, input/output token counts, error messages, solution correctness, and reasons for failure—in a pickle file to facilitate failure case analysis. The list of models integrated in npsolver is shown in Table 6.
Table 6: Online and offline models considered in this paper via npsolver.
Type Models Version Provider
Online GPT-4o-mini gpt-4o-mini-2024-07-18 OpenAI
GPT-4o gpt-4o-2024-08-06 OpenAI
o1-mini o1-mini-2024-09-12 OpenAI
o3-mini o3-mini-2025-01-31 OpenAI
DeepSeek-V3 deepseek-v3-241226 Huoshan
DeepSeek-V3-2503 deepseek-v3-250324 Huoshan
DeepSeek-R1 deepseek-r1-250120 Huoshan
Claude-3.7-Sonnet claude-3-7-sonnet-20250219 Anthropic
Offline QwQ-32B Qwen/QwQ-32B N/A
DeepSeek-R1-32B deepseek-ai/DeepSeek-R1-Distill-Qwen-32B N/A
Online.
The online state-of-the-art LLMs, e.g., o1/o3-mini and DeepSeek-v3/R1, can be accessed through APIs without local computational overhead. However, these online models have dependency on network stability and API costs with token usage. npsolver supports multiple providers, e.g., OpenAI, through modular API clients. We implement efficient batch processing with LiteLLM, which minimizes the latency during parallel problem-solving.
Offline.
Open-weight LLMs, e.g., QwQ-32B and Deepseek-R1-32B, can be accessed by deploying them locally. This allows for GPU-accelerated, high-throughput inference while avoiding API-related costs. Offline models are deployed using vLLM, with hyperparameters—such as temperature and maximum token length—manually configured according to their official technical documentation.
C.3 Evaluation Suite: npeval
npeval employs a statistically rigorous sampling strategy. For each difficulty, the aggregated performance over 3 different independent seeds, with 30 samples generated per seed, aligning with the minimum sample size for reliable statistical analysis [16], are considered. This sampling design, i.e., sampling 90 instances total per difficulty level for each problem, balances budget constraints while mitigating instance-specific variance.
Evaluation Metrics.
rliable [2] is an open-source Python library designed to enable statistically robust evaluation of reinforcement learning and machine learning benchmarks. Inspired by rliable, npeval provides the following 4 evaluation aggregate metrics:
  • Mean: Mean is a standard evaluation metric that treats each score equally and calculates the overall mean across runs and tasks.
  • Interquartile Mean (IQM): IQM trims extreme values and computes the interquartile mean across runs and tasks to smooth out the randomness in responses. IQM highlights the consistency of the performance and complements metrics like mean/median to avoid outlier skew.
  • Median: Median represents the middle value of the scores by calculating the median of the average scores per task across all runs, which is unaffected by extreme values.
  • Optimality Gap (OG): OG measures the average shortfall of scores below a predefined threshold $\gamma$, where all scores above $\gamma$ are clipped to $\gamma$, so as to quantify and penalize the underperformance, making it less susceptible to outliers compared to mean scores.
To quantify uncertainty in aggregate metrics, e.g. IQM, npeval employs stratified bootstrap confidence intervals (SBCIs) [11, 12] for the performance interval estimation. SBCIs use stratified resampling within predefined strata, e.g., difficulty levels, to preserve the hierarchical structure of the evaluation data, reduce bias, and provide statistically sound interval estimates.
Comprehensive Analysis
Based on evaluation metrics, npeval provides a comprehensive analysis of the LLMs’ performance over the problems and difficulty levels, including the full results for each problem, each model and each level (Appendix G), the performance over different problems (Appendix H), the analysis of both prompt and completion tokens of LLMs (Appendix I), the analysis of the number of “aha moments” during the DeepSeek-R1 reasoning [14] (Appendix J), an illustration of errors over problems (Table 5) with detailed error analysis (Appendix K), considering both the solution errors, i.e., the errors returned by npgym, and the reasoning errors, i.e., the errors produced in the internal reasoning process of LLMs, which enables the identification of the failure cases (Appendix L). The total cost of evaluation over different models is in listed in Table 23 (Appendix M).
Appendix D Prompts and Responses
Prompts.
In this section, we carefully design the prompt template of NPPC for LLMs to be simple, general, and consistent across different problems. The prompt template includes:
  • Problem description: where a concise definition of the NPC problem is provided, including the problem name, the input, and the question to be solved.
  • Examples: where one or multiple in context examples, defined as problem-solution pairs, are listed, demonstrating the expected solutions, i.e., answer correctness and format, for specific instances. These examples guide LLMs to generate the responses with the required format.
  • Problem to solve: a target instance that requires LLMs to generate the solution.
  • Instruction: which provides a directive to output answers in JSON format.
nppc_template = """
# <problem_name> Problem Description:
<problem_description>
# Examples:
<in_context_examples>
# Problem to Solve:
Problem: <problem_to_solve>
# Instruction:
Now please solve the above problem. Reason step by step and present your answer in the "solution" field in the following json format:
“‘json
{"solution": "___" }
example_and_solution = """Problem: <example_problem>
{"solution": <example_solution>}
Responses.
We extract the answers from the LLMs’ responses and the code is displayed below:
def extract_solution_from_response(response):
# find the json code
[…]
if not match:
[…]
if not match:
[…]
if match:
json_str = match[-1]
try:
# remove the single line comment
json_str = re.sub(r"//.*\$", "", json_str, flags=re.MULTILINE)
# remove the multiple line comment
[…]
data = json.loads(json_str)
answer = data["solution"]
return answer, None
except (json.JSONDecodeError, KeyError, SyntaxError) as e:
print(f"Error parsing JSON or answer field: {e}")
return None, f"Error parsing JSON or answer field: {e}"
else:
print("No JSON found in the text.")
return None, "JSON Error: No JSON found in the text."
The code extracts JSON data from LLM responses using three regex patterns in sequence:
  • First tries to find content between triple quotes with "json" marker,
  • If that fails, looks for "json" followed by content in curly braces,
  • If both fail, simply looks for any content between curly braces.
If all the three tries cannot find the content, we will raise the error.
Appendix E List of NP-complete Problems
Problem 1.
  • Name: 3-Satisfiability (3SAT)
  • Input: A set of $m$ clauses $\{C_{1},C_{2},\dots,C_{m}\}$ over a set of $n$ Boolean valued variables $X_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, such that each clause depends on exactly three distinct variables from $X_{n}$. A clause being a Boolean expression of the form $y_{i}\land y_{j}\land y_{k}$ where each $y$ is of the form $x$ or $\neg x$ (i.e. negation of $x$) with $x$ being some variable in $X_{n}$. For example if $n=4$ and $m=3$, a possible instance could be the (set of) Boolean expressions: $C_{1}=(x_{1}\land(\neg x_{2})\land(\neg x_{3}))$, $C_{2}=(x_{2}\land x_{3}\land(\neg x_{4}))$, $C_{3}=((\neg x_{1})\land x_{3}\land x_{4})$.
  • Question: Can each variable $x_{i}$ of $X_{n}$ be assigned a Boolean value $\alpha_{i}\in\{\text{true},\text{false}\}$ in such a way that every clause evaluates to the Boolean result true under the assignment $\langle x_{i}:=\alpha_{i},i\in\{1,\dots,n\}\rangle$?
Problem 2.
  • Name: Graph 3-Colourability (3-COL)
  • Input: An $n$-node undirected graph $G=(V,E)$ with node set $V$ and edge set $E$.
  • Question: Can each node of $G=(V,E)$ be assigned exactly one of three colours - Red, Blue, Green - in such a way that no two nodes which are joined by an edge, are assigned the same colour?
Problem 3.
  • Name: Clique
  • Input: An $n$-node undirected graph $G=(V,E)$ with node set $V$ and edge set $E$; a positive integer $k$ with $k\leq n$.
  • Question: Does $G$ contain a $k$-clique, i.e. a subset $W$ of the nodes $V$ such that $W$ has size $k$ and for each distinct pair of nodes $u,v$ in $W$, $\{u,v\}$ is an edge of $G$?
Problem 4.
  • Name: Vertex Cover
  • Input: An $n$-node undirected graph $G=(V,E)$ with node set $V$ and edge set $E$; a positive integer $k$ with $k\leq n$.
  • Question: Is there a subset $W$ of $V$ having size at most $k$ and such that for every edge $\{u,v\}$ in $E$ at least one of $u$ and $v$ belongs to $W$?
Problem 5.
  • Name: Quadratic Diophantine Equations
  • Input: Positive integers $a$, $b$, and $c$.
  • Question: Are there two positive integers $x$ and $y$ such that $(a*x*x)+(b*y)=c$?
Problem 6.
  • Name: Shortest Common Superstring
  • Input: A finite set $R=\{r_{1},r_{2},\dots,r_{m}\}$ of strings (sequences of symbols); positive integer $k$.
  • Question: Is there a string $w$ of length at most $k$ such that every string in $R$ is a substring of $w$, i.e., for each $r$ in $R$, $w$ can be decomposed as $w=w_{0}rw_{1}$ where $w_{0}$, $w_{1}$ are (possibly empty) strings?
Problem 7.
  • Name: Bandwidth
  • Input: $n$-node undirected graph $G=(V,E)$; positive integer $k\leq n$.
  • Question: Is there a linear ordering of $V$ with bandwidth at most $k$, i.e., a one-to-one function $f:V\rightarrow\{0,1,2,...,n-1\}$ such that for all edges ${u,v}$ in $G$, $|f(u)-f(v)|\leq k$?
Problem 8.
  • Name: Maximum Leaf Spanning Tree
  • Input: $n$-node undirected graph $G=(V,E)$; positive integer $k\leq n$.
  • Question: Does $G$ have a spanning tree in which at least $k$ nodes have degree 1?
Problem 9.
  • Name: Independent Set
  • Input: $n$-node undirected graph $G=(V,E)$; positive integer $k\leq n$.
  • Question: Does $G$ have an independent set of size at least $k$, i.e., a subset $W$ of at least $k$ nodes from $V$ such that no pair of nodes in $W$ is joined by an edge in $E$?
Problem 10.
  • Name: Hamiltonian Cycle
  • Input: n-node graph $G=(V,E)$.
  • Question: Is there a cycle in $G$ that visits every node in $V$ exactly once and returns to the starting node, and thus contains exactly $n$ edge
Problem 11.
  • Name: Travelling Salesman
[…]
  • Question: Is there an ordering $\langle\pi(1),\pi(2),...,\pi(n)\rangle$ of the $n$ cities such that the total travel distance, calculated as the sum of $d(\pi(i),\pi(i+1))$ for $i=1$ to $n-1$, plus $d(\pi(n),\pi(1))$, is at most $B$?
Problem 12.
  • Name: Dominating Set
  • Input: An undirected graph $G(V,E)$ with n nodes; a positive integer $k$ where $k\leq n$.
  • Question: Does $G$ contain a dominating set of size at most $k$, i.e. a subset $W$ of $V$ containing at most $k$ nodes such that every node $u$ in $V-W$ (i.e. in $V$ but not in $W$) has at least one neighbor $w$ in $W$ where ${u,w}$ is an edge in $E$?
Problem 13.
  • Name: 3-Dimensional Matching (3DM)
  • Input: 3 disjoint sets $X$, $Y$, and $Z$, each containing exactly $n$ elements; a set $M$ of $m$ triples $\{(x_{i},y_{i},z_{i}):1\leq i\leq m\}$ such that $x_{i}$ is in $X$, $y_{i}$ in $Y$, and $z_{i}$ in $Z$, i.e. $M$ is a subset of $X\times Y\times Z$.
  • Question: Does $M$ contain a matching, i.e., is there a subset $Q$ of $M$ such that $|Q|=n$ and for all distinct pairs of triples $(u,v,w)$ and $(x,y,z)$ in $Q$ it holds that $u\neq x$ and $v\neq y$ and $w\neq z$?
Problem 14.
  • Name: Set Splitting
  • Input: A finite set $S$; A collection $C_{1},\dots,C_{m}$ of subsets of $S$.
  • Question: Can $S$ be partitioned into two disjoint subsets - $S1$ and $S2$ such that for each set $C_{i}$ it holds that $C_{i}$ is not a subset of $S_{1}$ and $C_{i}$ is not a subset of $S_{2}$?
Problem 15.
  • Name: Set Packing
  • Input: A collection $C=(C_{1},\dots,C_{m})$ of finite sets; a positive integer $k\leq m$.
  • Question: Are there $k$ sets - $D_{1},\dots,D_{k}$ from the collection $C$ such that for all $1\leq i<j\leq k$, $D_{i}$ and $D_{j}$ have no common elements?
Problem 16.
  • Name: Exact Cover by 3-Sets (X3C)
  • Input: A finite set $X$ containing exactly $3n$ elements; a collection $C$ of subsets of $X$ each of which contains exactly $3$ elements.
  • Question: Does $C$ contain an exact cover for $X$, i.e., a sub-collection of 3-element sets $D=(D_{1},\dots,D_{n})$ such that each element of $X$ occurs in exactly one subset in $D$?
Problem 17.
  • Name: Minimum Cover
  • Input: A finite set $S$; A collection $C=(C_{1},\dots,C_{m})$ of subsets of $S$; a positive integer $k\leq m$.
  • Question: Does $C$ contain a cover for $S$ comprising at most $k$ subsets, i.e., a collection $D=(D_{1},\dots,D_{t})$, where $t\leq k$, each $D_{i}$ is a set in $C$, and such that every element in $S$ belongs to at least one set in $D$?
Problem 18.
  • Name: Partition
  • Input: Finite set $A$; for each element $a$ in $A$ a positive integer size $s(a)$.
  • Question: Can $A$ be partitioned into $2$ disjoint sets $A_{1}$ and $A_{2}$ in a such a way that $\sum_{a\in A_{1}}s(a)=\sum_{a\in A_{2}}s(a)$?
Problem 19.
  • Name: Subset Sum
  • Input: Finite set $A$; for each element $a\in A$ a positive integer size $s(a)$; a positive integer $K$.
  • Question: Is there a subset $B$ of $A$ such that $\sum_{a\in B}s(a)=K$?
Problem 20.
  • Name: Minimum Sum of Squares
  • Input: A set $A$ of $n$ elements; for each element $a\in A$ a positive integer size $s(a)$; positive integers $k\leq n$ and $J$.
  • Question: Can $A$ be partitioned into $k$ disjoint sets $A_{1},\dots,A_{k}$ such that $\sum_{i=1}^{k}(\sum_{x\in A_{i}}s(x))^{2}<=J$?
Problem 21.
  • Name: Bin Packing
  • Input: A finite set $U$ of $m$ items; for each item $u$ in $U$ a positive integer size $s(u)$; positive integers $B$ (bin capacity) and $k$, where $k\leq m$.
  • Question: Can $U$ be partitioned into $k$ disjoint sets $U_{1},\dots,U_{k}$ such that the total size of the items in each subset $U_{i}$ (for $1\leq i\leq k$) does not exceed $B$?
Problem 22.
  • Name: Hitting String
  • Input: Finite set $S=\{s_{1},\dots,s_{m}\}$ each $s_{i}$ being a string of $n$ symbols over $\{0,1,*\}$.
  • Question: Is there a binary string $x=x_{1}x_{2}\dots x_{n}$ of length $n$ such that for each $s_{j}\in S$, $s_{j}$ and $x$ agree in at least one position?
Problem 23.
  • Name: Quadratic Congruences
  • Input: Positive integers $a$, $b$, and $c$.
  • Question: Is there a positive integer $x$ whose value is less than $c$ and is such that $x^{2}\mod b==a$, i.e., the remainder when $x^{2}$ is divided by $b$ is equal to $a$?
Problem 24.
  • Name: Betweenness
  • Input: A finite set $A$ of size n; a set $C$ of ordered triples, $(a,b,c)$, of distinct elements from $A$.
  • Question: Is there a one-to-one function, $f:A\rightarrow\{0,1,2,...,n-1\}$ such that for each triple $(a,b,c)$ in $C$ it holds that either $f(a)<f(b)<f(c)$ or $f(c)<f(b)<f(a)$?
Problem 25.
  • Name: Clustering
  • Input: Finite set $X$; for each pair of elements $x$ and $y$ in $X$, a positive integer distance $d(x,y)$; positive integer $B$.
  • Question: Is there a partition of $X$ into 3 disjoint sets - $X_{1},X_{2},X_{3}$ with which: for each set $X_{i},i\in\{1,2,3\}$, for all pairs $x$ and $y$ in $X_{i}$, it holds that $d(x,y)\leq B$?
Appendix F Hyperparameters
The hyperparameters used for benchmarking are listed in Table 7. For both offline and online-deployed models, accuracy is averaged over three seeds and 30 trials per difficulty level per task. Each model is allowed up to three attempts to mitigate the impact of API connection issues. For offline models, we follow the recommended sampling parameters from the technical reports of Deepseek-R1-32B and QwQ-32B for vLLM deployment.
Table 7: Hyperparameters
Type Hyperparameter Value
Basic seeds 42, 53, 64
n_shots 1
n_trials 30
batch_size 10
max_tries 3
Offline Model temperature 0.6
top_p 0.95
max_tokens 7500
gpu_memory_utilization 0.8
Appendix G Full Results over Problems
In this section, we present the full results over problems, as displayed in Figure 8. For each element in the table $x^{a}_{b}$, $x$ is the value of IQM and $a$ and $b$ are the upper and lower values of the CI, respectively.
Table 8: 3SAT
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.94_{0.90}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.56_{0.53}^{0.60}$ $0.11_{0.07}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.83_{0.73}^{0.90}$ $0.52_{0.40}^{0.70}$ $0.32_{0.30}^{0.37}$ $0.19_{0.13}^{0.27}$ $0.13_{0.07}^{0.23}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.84_{0.83}^{0.87}$ $0.27_{0.20}^{0.30}$ $0.17_{0.10}^{0.27}$ $0.08_{0.03}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.94_{0.93}^{0.97}$ $0.51_{0.47}^{0.57}$ $0.43_{0.40}^{0.47}$ $0.22_{0.20}^{0.27}$ $0.09_{0.00}^{0.17}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.89_{0.80}^{0.93}$ $0.62_{0.60}^{0.67}$ $0.54_{0.50}^{0.60}$ $0.36_{0.27}^{0.47}$ $0.19_{0.13}^{0.27}$ $0.14_{0.03}^{0.23}$ $0.08_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3 $0.94_{0.93}^{0.97}$ $0.78_{0.60}^{0.90}$ $0.38_{0.33}^{0.40}$ $0.34_{0.17}^{0.43}$ $0.21_{0.17}^{0.27}$ $0.06_{0.03}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.97}$ $0.68_{0.60}^{0.80}$ $0.53_{0.47}^{0.63}$ $0.38_{0.30}^{0.43}$ $0.28_{0.23}^{0.33}$ $0.12_{0.03}^{0.23}$ $0.08_{0.03}^{0.17}$ $0.03_{0.03}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.98_{0.93}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.83_{0.63}^{0.93}$ $0.64_{0.63}^{0.67}$ $0.23_{0.20}^{0.27}$ $0.13_{0.10}^{0.17}$
o1-mini $0.92_{0.90}^{0.93}$ $0.91_{0.87}^{0.97}$ $0.92_{0.90}^{0.97}$ $0.81_{0.77}^{0.87}$ $0.67_{0.60}^{0.77}$ $0.20_{0.10}^{0.37}$ $0.03_{0.03}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.93_{0.90}^{0.97}$ $0.82_{0.77}^{0.87}$ $0.72_{0.63}^{0.83}$ $0.77_{0.70}^{0.83}$ $0.82_{0.80}^{0.83}$ $0.71_{0.67}^{0.77}$ $0.60_{0.53}^{0.70}$ $0.30_{0.20}^{0.43}$ $0.13_{0.10}^{0.17}$ $0.12_{0.03}^{0.17}$
Table 9: Vertex Cover
1 2 3 4 5 6 7 8 9 10
QwQ-32B $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.50_{0.37}^{0.60}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.91_{0.83}^{1.00}$ $0.92_{0.90}^{0.93}$ $0.81_{0.73}^{0.87}$ $0.52_{0.43}^{0.60}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.94_{0.87}^{1.00}$ $0.67_{0.57}^{0.80}$ $0.37_{0.27}^{0.43}$ $0.18_{0.10}^{0.23}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.96_{0.90}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $0.60_{0.57}^{0.63}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.97_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.53_{0.47}^{0.57}$ $0.37_{0.30}^{0.47}$ $0.37_{0.23}^{0.50}$ $0.26_{0.20}^{0.30}$ $0.14_{0.10}^{0.17}$ $0.04_{0.00}^{0.07}$
DeepSeek-V3 $0.92_{0.87}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.34_{0.23}^{0.43}$ $0.14_{0.10}^{0.20}$ $0.06_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.28_{0.10}^{0.43}$ $0.37_{0.23}^{0.50}$ $0.27_{0.23}^{0.33}$ $0.09_{0.07}^{0.13}$ $0.09_{0.07}^{0.10}$ $0.01_{0.00}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.77_{0.70}^{0.87}$ $0.41_{0.33}^{0.47}$ $0.18_{0.13}^{0.20}$ $0.13_{0.08}^{0.20}$ $0.06_{0.00}^{0.10}$
o1-mini $0.74_{0.73}^{0.77}$ $0.77_{0.73}^{0.80}$ $0.78_{0.70}^{0.83}$ $0.91_{0.87}^{0.93}$ $0.58_{0.43}^{0.70}$ $0.31_{0.27}^{0.33}$ $0.13_{0.10}^{0.17}$ $0.13_{0.03}^{0.27}$ $0.08_{0.07}^{0.10}$ $0.02_{0.00}^{0.07}$
o3-mini $0.82_{0.70}^{0.90}$ $0.89_{0.83}^{0.93}$ $0.89_{0.83}^{0.93}$ $0.80_{0.73}^{0.90}$ $0.59_{0.53}^{0.70}$ $0.52_{0.50}^{0.57}$ $0.19_{0.10}^{0.27}$ $0.13_{0.03}^{0.23}$ $0.11_{0.03}^{0.17}$ $0.07_{0.03}^{0.10}$
Table 10: Superstring
1 2 3 4 5 6 7 8 9 10
QwQ-32B $1.00_{1.00}^{1.00}$ $0.92_{0.87}^{0.97}$ $0.28_{0.20}^{0.33}$ $0.19_{0.13}^{0.23}$ $0.17_{0.10}^{0.23}$ $0.06_{0.00}^{0.13}$ $0.08_{0.03}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.58_{0.33}^{0.70}$ $0.24_{0.10}^{0.40}$ $0.16_{0.07}^{0.23}$ $0.12_{0.07}^{0.23}$ $0.10_{0.07}^{0.17}$ $0.03_{0.03}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.32_{0.20}^{0.47}$ $0.08_{0.03}^{0.13}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.81_{0.77}^{0.83}$ $0.47_{0.30}^{0.57}$ $0.11_{0.07}^{0.17}$ $0.10_{0.03}^{0.17}$ $0.06_{0.03}^{0.10}$ $0.16_{0.03}^{0.27}$ $0.06_{0.03}^{0.10}$ $0.12_{0.10}^{0.13}$ $0.07_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.78_{0.70}^{0.90}$ $0.51_{0.47}^{0.57}$ $0.68_{0.60}^{0.77}$ $0.74_{0.70}^{0.80}$ $0.77_{0.73}^{0.80}$ $0.88_{0.83}^{0.93}$ $0.82_{0.77}^{0.90}$ $0.74_{0.70}^{0.80}$
DeepSeek-V3 $0.80_{0.73}^{0.83}$ $0.52_{0.37}^{0.67}$ $0.49_{0.43}^{0.53}$ $0.46_{0.40}^{0.53}$ $0.44_{0.40}^{0.53}$ $0.40_{0.30}^{0.57}$ $0.24_{0.13}^{0.37}$ $0.22_{0.17}^{0.27}$ $0.08_{0.03}^{0.10}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.99_{0.97}^{1.00}$ $0.89_{0.83}^{0.93}$ $0.78_{0.73}^{0.83}$ $0.61_{0.57}^{0.67}$ $0.53_{0.40}^{0.60}$ $0.37_{0.33}^{0.40}$ $0.21_{0.20}^{0.23}$ $0.17_{0.17}^{0.17}$ $0.26_{0.23}^{0.27}$ $0.13_{0.10}^{0.17}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.81_{0.73}^{0.90}$ $0.80_{0.73}^{0.83}$ $0.61_{0.53}^{0.77}$ $0.37_{0.20}^{0.50}$ $0.31_{0.30}^{0.33}$ $0.11_{0.07}^{0.17}$ $0.13_{0.10}^{0.17}$
o1-mini $0.91_{0.87}^{0.97}$ $0.59_{0.47}^{0.73}$ $0.48_{0.43}^{0.53}$ $0.20_{0.17}^{0.23}$ $0.10_{0.07}^{0.13}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.89_{0.87}^{0.90}$ $0.74_{0.70}^{0.77}$ $0.31_{0.23}^{0.37}$ $0.04_{0.03}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Table 11: QDE
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.72_{0.30}^{1.00}$ $0.56_{0.17}^{0.80}$ $0.19_{0.07}^{0.27}$ $0.16_{0.07}^{0.23}$ $0.03_{0.03}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.84_{0.77}^{0.90}$ $0.62_{0.50}^{0.70}$ $0.11_{0.10}^{0.13}$ $0.08_{0.07}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.49_{0.43}^{0.53}$ $0.23_{0.17}^{0.27}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.67_{0.60}^{0.70}$ $0.43_{0.33}^{0.57}$ $0.08_{0.07}^{0.10}$ $0.03_{0.03}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.87}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.78_{0.77}^{0.80}$ $0.59_{0.47}^{0.67}$ $0.10_{0.07}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.97_{0.97}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.38_{0.37}^{0.40}$ $0.19_{0.10}^{0.30}$ $0.04_{0.03}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.68_{0.63}^{0.70}$ $0.64_{0.57}^{0.73}$ $0.30_{0.20}^{0.37}$ $0.17_{0.13}^{0.20}$ $0.08_{0.00}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.08_{0.00}^{0.13}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.68_{0.63}^{0.73}$ $0.27_{0.20}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.09_{0.07}^{0.13}$ $0.03_{0.00}^{0.07}$
o1-mini $0.57_{0.50}^{0.70}$ $0.59_{0.50}^{0.63}$ $0.44_{0.33}^{0.63}$ $0.46_{0.43}^{0.50}$ $0.11_{0.07}^{0.17}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.94_{0.90}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.96_{0.90}^{1.00}$ $0.81_{0.77}^{0.87}$ $0.66_{0.53}^{0.77}$ $0.30_{0.20}^{0.43}$ $0.27_{0.20}^{0.30}$ $0.27_{0.23}^{0.30}$ $0.13_{0.10}^{0.17}$
Table 12: 3DM
1 2 3 4 5 6 7 8 9 10
QwQ-32B $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.94_{0.93}^{0.97}$ $0.33_{0.07}^{0.83}$ $0.06_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.87_{0.77}^{1.00}$ $0.42_{0.23}^{0.57}$ $0.09_{0.03}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.43_{0.27}^{0.57}$ $0.09_{0.07}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.64_{0.53}^{0.83}$ $0.24_{0.17}^{0.37}$ $0.13_{0.03}^{0.20}$ $0.10_{0.07}^{0.13}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.93}^{0.97}$ $0.84_{0.80}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.59_{0.43}^{0.70}$ $0.21_{0.13}^{0.33}$ $0.09_{0.07}^{0.10}$ $0.07_{0.03}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.74_{0.57}^{0.83}$ $0.32_{0.23}^{0.47}$ $0.08_{0.00}^{0.13}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.94_{0.93}^{0.97}$ $0.76_{0.70}^{0.87}$ $0.49_{0.43}^{0.60}$ $0.31_{0.10}^{0.47}$ $0.07_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.91_{0.83}^{0.97}$ $0.91_{0.87}^{0.97}$ $0.57_{0.50}^{0.67}$ $0.27_{0.17}^{0.37}$ $0.02_{0.00}^{0.03}$
o1-mini $0.87_{0.83}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.81_{0.73}^{0.87}$ $0.77_{0.73}^{0.83}$ $0.38_{0.30}^{0.47}$ $0.26_{0.23}^{0.27}$ $0.11_{0.07}^{0.20}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.63_{0.60}^{0.70}$ $0.86_{0.80}^{0.93}$ $0.72_{0.67}^{0.77}$ $0.71_{0.57}^{0.80}$ $0.57_{0.53}^{0.60}$ $0.56_{0.43}^{0.70}$ $0.38_{0.30}^{0.53}$ $0.30_{0.23}^{0.37}$ $0.23_{0.23}^{0.23}$ $0.20_{0.17}^{0.23}$
Table 13: TSP
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.61_{0.50}^{0.80}$ $0.41_{0.30}^{0.53}$ $0.42_{0.30}^{0.53}$ $0.56_{0.50}^{0.60}$ $0.26_{0.23}^{0.30}$ $0.19_{0.13}^{0.27}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.88_{0.87}^{0.90}$ $0.62_{0.53}^{0.73}$ $0.30_{0.23}^{0.40}$ $0.13_{0.03}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.93_{0.90}^{0.96}$ $0.34_{0.27}^{0.40}$ $0.12_{0.07}^{0.20}$ $0.07_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.97_{0.97}^{0.97}$ $0.76_{0.73}^{0.80}$ $0.59_{0.50}^{0.67}$ $0.40_{0.33}^{0.47}$ $0.22_{0.10}^{0.33}$ $0.16_{0.03}^{0.23}$ $0.08_{0.07}^{0.10}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.83_{0.77}^{0.90}$ $0.86_{0.80}^{0.90}$ $0.80_{0.73}^{0.83}$ $0.54_{0.47}^{0.70}$ $0.51_{0.50}^{0.53}$ $0.08_{0.03}^{0.10}$ $0.06_{0.00}^{0.10}$
DeepSeek-V3 $0.98_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.74_{0.60}^{0.83}$ $0.62_{0.50}^{0.77}$ $0.49_{0.33}^{0.67}$ $0.49_{0.33}^{0.73}$ $0.17_{0.13}^{0.23}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{0.97}$ $0.96_{0.87}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.70_{0.67}^{0.77}$ $0.66_{0.57}^{0.70}$ $0.39_{0.27}^{0.47}$ $0.10_{0.10}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.87_{0.87}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.62_{0.57}^{0.67}$ $0.24_{0.20}^{0.30}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$
o1-mini $0.84_{0.73}^{0.90}$ $0.89_{0.87}^{0.93}$ $0.67_{0.60}^{0.77}$ $0.57_{0.43}^{0.63}$ $0.34_{0.23}^{0.47}$ $0.37_{0.23}^{0.43}$ $0.18_{0.07}^{0.30}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.79_{0.73}^{0.87}$ $0.62_{0.53}^{0.67}$ $0.53_{0.47}^{0.63}$ $0.28_{0.20}^{0.33}$ $0.31_{0.23}^{0.37}$ $0.30_{0.17}^{0.47}$ $0.30_{0.20}^{0.37}$ $0.19_{0.17}^{0.20}$ $0.12_{0.07}^{0.17}$ $0.07_{0.00}^{0.13}$
Table 14: Hamiltonian Cycle
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.94_{0.90}^{1.00}$ $0.87_{0.83}^{0.93}$ $0.80_{0.70}^{0.93}$ $0.62_{0.57}^{0.67}$ $0.33_{0.23}^{0.40}$ $0.16_{0.10}^{0.20}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.69_{0.67}^{0.73}$ $0.36_{0.27}^{0.40}$ $0.24_{0.13}^{0.40}$ $0.09_{0.03}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.70_{0.67}^{0.73}$ $0.26_{0.13}^{0.40}$ $0.09_{0.07}^{0.10}$ $0.08_{0.03}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.73_{0.73}^{0.73}$ $0.39_{0.33}^{0.43}$ $0.22_{0.17}^{0.27}$ $0.12_{0.07}^{0.20}$ $0.09_{0.00}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.06_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.80_{0.70}^{0.90}$ $0.74_{0.67}^{0.83}$ $0.64_{0.53}^{0.77}$ $0.32_{0.17}^{0.50}$ $0.23_{0.20}^{0.27}$ $0.27_{0.20}^{0.33}$ $0.16_{0.10}^{0.27}$ $0.10_{0.10}^{0.10}$ $0.02_{0.00}^{0.07}$
DeepSeek-V3 $0.83_{0.77}^{0.90}$ $0.44_{0.30}^{0.53}$ $0.14_{0.10}^{0.20}$ $0.16_{0.13}^{0.17}$ $0.09_{0.00}^{0.17}$ $0.06_{0.03}^{0.07}$ $0.06_{0.00}^{0.10}$ $0.06_{0.00}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.99_{0.97}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.51_{0.50}^{0.53}$ $0.38_{0.27}^{0.53}$ $0.16_{0.07}^{0.27}$ $0.14_{0.13}^{0.17}$ $0.09_{0.07}^{0.10}$ $0.10_{0.07}^{0.17}$ $0.06_{0.03}^{0.07}$ $0.03_{0.00}^{0.07}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.83}^{0.97}$ $0.76_{0.63}^{0.93}$ $0.64_{0.57}^{0.73}$ $0.49_{0.43}^{0.57}$ $0.36_{0.27}^{0.40}$ $0.17_{0.13}^{0.23}$ $0.04_{0.00}^{0.10}$
o1-mini $0.72_{0.57}^{0.80}$ $0.71_{0.67}^{0.77}$ $0.54_{0.50}^{0.60}$ $0.40_{0.33}^{0.47}$ $0.19_{0.17}^{0.23}$ $0.23_{0.13}^{0.30}$ $0.12_{0.10}^{0.17}$ $0.08_{0.03}^{0.10}$ $0.04_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$
o3-mini $0.82_{0.80}^{0.83}$ $0.84_{0.77}^{0.90}$ $0.71_{0.63}^{0.77}$ $0.71_{0.60}^{0.83}$ $0.63_{0.57}^{0.73}$ $0.59_{0.50}^{0.67}$ $0.44_{0.37}^{0.50}$ $0.32_{0.27}^{0.43}$ $0.20_{0.10}^{0.33}$ $0.22_{0.20}^{0.23}$
Table 15: Bin Packing
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.88_{0.80}^{0.93}$ $0.83_{0.77}^{0.87}$ $0.46_{0.40}^{0.57}$ $0.08_{0.00}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.26_{0.17}^{0.33}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.30_{0.23}^{0.33}$ $0.03_{0.03}^{0.03}$ $0.04_{0.00}^{0.10}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.83_{0.73}^{0.90}$ $0.44_{0.40}^{0.50}$ $0.34_{0.33}^{0.37}$ $0.18_{0.13}^{0.20}$ $0.04_{0.00}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.93}$ $0.58_{0.43}^{0.70}$ $0.39_{0.37}^{0.40}$ $0.07_{0.00}^{0.17}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.66_{0.60}^{0.73}$ $0.46_{0.40}^{0.50}$ $0.44_{0.37}^{0.50}$ $0.37_{0.33}^{0.40}$ $0.06_{0.00}^{0.13}$ $0.04_{0.03}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.87_{0.80}^{0.93}$ $0.74_{0.67}^{0.83}$ $0.62_{0.57}^{0.67}$ $0.18_{0.10}^{0.27}$ $0.18_{0.13}^{0.23}$ $0.09_{0.03}^{0.13}$ $0.02_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.80_{0.73}^{0.90}$ $0.64_{0.57}^{0.77}$ $0.49_{0.47}^{0.53}$ $0.29_{0.20}^{0.43}$ $0.06_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$
o1-mini $0.67_{0.43}^{0.80}$ $0.58_{0.50}^{0.63}$ $0.52_{0.47}^{0.57}$ $0.33_{0.27}^{0.40}$ $0.31_{0.20}^{0.50}$ $0.19_{0.13}^{0.23}$ $0.07_{0.03}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.02_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
o3-mini $0.72_{0.60}^{0.83}$ $0.67_{0.57}^{0.80}$ $0.62_{0.57}^{0.67}$ $0.48_{0.33}^{0.57}$ $0.41_{0.37}^{0.47}$ $0.29_{0.17}^{0.43}$ $0.24_{0.13}^{0.47}$ $0.17_{0.10}^{0.27}$ $0.42_{0.33}^{0.47}$ $0.28_{0.20}^{0.33}$
Table 16: 3-COL
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.87}^{0.93}$ $0.78_{0.73}^{0.87}$ $0.56_{0.50}^{0.67}$ $0.34_{0.27}^{0.43}$ $0.10_{0.07}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.49_{0.43}^{0.57}$ $0.51_{0.43}^{0.57}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.40_{0.30}^{0.50}$ $0.17_{0.13}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.60_{0.57}^{0.63}$ $0.39_{0.30}^{0.53}$ $0.03_{0.00}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.76_{0.67}^{0.87}$ $0.70_{0.67}^{0.73}$ $0.22_{0.07}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.09_{0.07}^{0.10}$ $0.04_{0.03}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.67_{0.63}^{0.70}$ $0.60_{0.53}^{0.63}$ $0.13_{0.07}^{0.20}$ $0.12_{0.10}^{0.17}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.80_{0.73}^{0.87}$ $0.90_{0.87}^{0.93}$ $0.48_{0.40}^{0.63}$ $0.64_{0.57}^{0.73}$ $0.32_{0.30}^{0.37}$ $0.16_{0.07}^{0.20}$ $0.16_{0.07}^{0.23}$ $0.09_{0.00}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $0.99_{0.97}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.88_{0.80}^{0.93}$ $0.72_{0.67}^{0.77}$ $0.72_{0.67}^{0.80}$ $0.51_{0.40}^{0.67}$ $0.22_{0.17}^{0.27}$ $0.04_{0.03}^{0.07}$
o1-mini $0.61_{0.50}^{0.70}$ $0.76_{0.70}^{0.87}$ $0.57_{0.43}^{0.70}$ $0.62_{0.60}^{0.67}$ $0.37_{0.33}^{0.43}$ $0.27_{0.23}^{0.30}$ $0.34_{0.27}^{0.40}$ $0.17_{0.07}^{0.23}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$
o3-mini $0.98_{0.97}^{1.00}$ $0.91_{0.87}^{0.93}$ $0.96_{0.90}^{1.00}$ $0.84_{0.83}^{0.87}$ $0.78_{0.70}^{0.87}$ $0.72_{0.67}^{0.80}$ $0.71_{0.60}^{0.80}$ $0.61_{0.47}^{0.80}$ $0.51_{0.47}^{0.53}$ $0.29_{0.27}^{0.30}$
Table 17: Min Sum Square
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.77_{0.70}^{0.80}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.23_{0.10}^{0.43}$ $0.00_{0.00}^{0.00}$ $0.04_{0.00}^{0.07}$ $0.07_{0.03}^{0.10}$ $0.06_{0.03}^{0.10}$ $0.04_{0.00}^{0.13}$ $0.02_{0.00}^{0.07}$ $0.06_{0.03}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.06_{0.03}^{0.07}$
GPT-4o-mini $0.74_{0.67}^{0.80}$ $0.62_{0.50}^{0.77}$ $0.03_{0.00}^{0.07}$ $0.07_{0.03}^{0.10}$ $0.08_{0.07}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.04_{0.00}^{0.13}$
GPT-4o $0.94_{0.90}^{0.97}$ $0.82_{0.80}^{0.87}$ $0.46_{0.37}^{0.53}$ $0.56_{0.53}^{0.60}$ $0.44_{0.40}^{0.53}$ $0.48_{0.43}^{0.53}$ $0.04_{0.03}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
Claude-3.7-Sonnet $0.98_{0.97}^{1.00}$ $0.84_{0.80}^{0.93}$ $0.83_{0.80}^{0.90}$ $0.73_{0.70}^{0.80}$ $0.79_{0.70}^{0.87}$ $0.64_{0.60}^{0.70}$ $0.59_{0.50}^{0.63}$ $0.67_{0.57}^{0.73}$ $0.62_{0.53}^{0.67}$ $0.14_{0.07}^{0.20}$
DeepSeek-V3 $0.87_{0.80}^{0.93}$ $0.90_{0.83}^{0.93}$ $0.84_{0.80}^{0.90}$ $0.58_{0.53}^{0.63}$ $0.58_{0.53}^{0.63}$ $0.48_{0.43}^{0.57}$ $0.07_{0.03}^{0.10}$ $0.17_{0.07}^{0.23}$ $0.07_{0.00}^{0.17}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.48_{0.40}^{0.57}$ $0.71_{0.67}^{0.77}$ $0.59_{0.50}^{0.63}$ $0.62_{0.57}^{0.67}$ $0.61_{0.50}^{0.70}$ $0.22_{0.20}^{0.23}$ $0.29_{0.23}^{0.37}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.64_{0.57}^{0.73}$ $0.46_{0.37}^{0.53}$ $0.47_{0.33}^{0.53}$ $0.39_{0.30}^{0.43}$ $0.13_{0.10}^{0.17}$ $0.13_{0.10}^{0.17}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$
o1-mini $0.62_{0.57}^{0.67}$ $0.70_{0.63}^{0.80}$ $0.27_{0.23}^{0.30}$ $0.18_{0.10}^{0.27}$ $0.14_{0.10}^{0.23}$ $0.10_{0.07}^{0.13}$ $0.03_{0.00}^{0.10}$ $0.06_{0.03}^{0.07}$ $0.02_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
o3-mini $0.69_{0.60}^{0.80}$ $0.38_{0.23}^{0.47}$ $0.38_{0.33}^{0.47}$ $0.39_{0.23}^{0.50}$ $0.52_{0.47}^{0.60}$ $0.30_{0.23}^{0.37}$ $0.44_{0.33}^{0.50}$ $0.24_{0.20}^{0.27}$ $0.03_{0.00}^{0.07}$ $0.18_{0.13}^{0.23}$
Table 18: Bandwidth
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.90_{0.83}^{1.00}$ $0.84_{0.70}^{0.93}$ $0.76_{0.60}^{0.87}$ $0.66_{0.60}^{0.77}$ $0.63_{0.60}^{0.67}$ $0.30_{0.20}^{0.40}$ $0.20_{0.13}^{0.30}$ $0.06_{0.03}^{0.07}$
DeepSeek-R1-32B $0.93_{0.83}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.87_{0.80}^{0.93}$ $0.67_{0.53}^{0.83}$ $0.54_{0.47}^{0.70}$ $0.49_{0.43}^{0.57}$ $0.38_{0.37}^{0.40}$ $0.11_{0.07}^{0.13}$ $0.14_{0.10}^{0.17}$ $0.03_{0.00}^{0.07}$
GPT-4o-mini $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.84_{0.83}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.47_{0.40}^{0.57}$ $0.46_{0.43}^{0.47}$ $0.20_{0.17}^{0.23}$ $0.14_{0.10}^{0.20}$ $0.03_{0.00}^{0.07}$
GPT-4o $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.94_{0.87}^{1.00}$ $0.78_{0.67}^{0.87}$ $0.62_{0.57}^{0.67}$ $0.60_{0.53}^{0.67}$ $0.22_{0.17}^{0.30}$ $0.10_{0.07}^{0.13}$ $0.02_{0.00}^{0.03}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $0.66_{0.60}^{0.73}$ $0.62_{0.57}^{0.67}$ $0.28_{0.17}^{0.33}$ $0.11_{0.07}^{0.13}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3 $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.74_{0.63}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.56_{0.50}^{0.63}$ $0.34_{0.30}^{0.40}$ $0.23_{0.17}^{0.33}$ $0.03_{0.00}^{0.07}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.62_{0.57}^{0.67}$ $0.58_{0.53}^{0.63}$ $0.57_{0.53}^{0.60}$ $0.43_{0.33}^{0.50}$ $0.33_{0.27}^{0.40}$ $0.17_{0.13}^{0.20}$ $0.04_{0.03}^{0.07}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.93_{0.90}^{1.00}$ $0.88_{0.80}^{0.93}$ $0.83_{0.80}^{0.90}$ $0.68_{0.57}^{0.77}$ $0.59_{0.47}^{0.67}$ $0.34_{0.30}^{0.43}$ $0.24_{0.20}^{0.30}$ $0.07_{0.03}^{0.10}$
o1-mini $0.74_{0.70}^{0.80}$ $0.74_{0.60}^{0.83}$ $0.84_{0.83}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.68_{0.67}^{0.70}$ $0.59_{0.53}^{0.63}$ $0.33_{0.30}^{0.37}$ $0.24_{0.20}^{0.33}$ $0.06_{0.03}^{0.07}$
o3-mini $0.80_{0.77}^{0.83}$ $0.88_{0.83}^{0.93}$ $0.82_{0.77}^{0.93}$ $0.90_{0.87}^{0.93}$ $0.72_{0.60}^{0.80}$ $0.58_{0.43}^{0.70}$ $0.52_{0.50}^{0.53}$ $0.20_{0.17}^{0.23}$ $0.17_{0.13}^{0.20}$ $0.08_{0.07}^{0.10}$
Table 19: Max Leaf Span Tree
1 2 3 4 5 6 7 8 9 10
QwQ-32B $0.73_{0.57}^{0.83}$ $0.93_{0.87}^{0.97}$ $0.28_{0.20}^{0.40}$ $0.06_{0.03}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.20_{0.03}^{0.30}$ $0.24_{0.13}^{0.37}$ $0.18_{0.13}^{0.27}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.26_{0.17}^{0.33}$ $0.19_{0.07}^{0.40}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.49_{0.40}^{0.57}$ $0.53_{0.47}^{0.60}$ $0.29_{0.23}^{0.37}$ $0.24_{0.20}^{0.30}$ $0.08_{0.07}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.82_{0.70}^{0.93}$ $0.71_{0.57}^{0.83}$ $0.59_{0.57}^{0.63}$ $0.12_{0.07}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.79_{0.77}^{0.83}$ $0.88_{0.80}^{0.93}$ $0.89_{0.80}^{0.93}$ $0.69_{0.57}^{0.80}$ $0.56_{0.50}^{0.60}$ $0.26_{0.20}^{0.33}$ $0.27_{0.13}^{0.43}$ $0.09_{0.00}^{0.17}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.90_{0.80}^{0.97}$ $0.86_{0.83}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.39_{0.33}^{0.43}$ $0.18_{0.10}^{0.30}$ $0.22_{0.13}^{0.27}$ $0.28_{0.23}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$
DeepSeek-R1 $0.97_{0.97}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.88_{0.87}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.39_{0.37}^{0.43}$ $0.18_{0.13}^{0.23}$ $0.21_{0.20}^{0.23}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
o1-mini $0.70_{0.67}^{0.73}$ $0.53_{0.53}^{0.53}$ $0.57_{0.57}^{0.57}$ $0.17_{0.10}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
o3-mini $0.77_{0.70}^{0.83}$ $0.68_{0.63}^{0.73}$ $0.66_{0.63}^{0.67}$ $0.66_{0.60}^{0.70}$ $0.42_{0.30}^{0.50}$ $0.26_{0.23}^{0.27}$ $0.19_{0.17}^{0.23}$ $0.16_{0.07}^{0.33}$ $0.11_{0.03}^{0.17}$ $0.09_{0.03}^{0.17}$
Appendix H Performance over Problems
In this section, we present the performance of LLMs on each problem across different levels.
Figure 15: 3SAT
Figure 16: Vertex Cover
Figure 17: Superstring
Figure 18: QDE
Figure 19: 3DM
Figure 20: TSP
Figure 21: Hamiltonian Cycle
Figure 22: Bin Packing
Figure 23: 3-COL
Figure 24: Min Sum Square
Figure 25: Max Leaf Span Tree
Figure 26: Bandwidth
Appendix I Tokens
In this section, we present the results of the prompt and completion tokens used in LLMs.
Figure 27: 3SAT
Figure 28: Vertex Cover
Figure 29: Superstring
Figure 30: QDE
Figure 31: 3DM
Figure 32: TSP
Figure 33: Hamiltonian Cycle
Figure 34: Bin Packing
Figure 35: 3-COL
Figure 36: Min Sum Square
Figure 37: Max Leaf Span Tree
Figure 38: Bandwidth
Appendix J Aha Moments
This section investigates the phenomenon of “aha moments”, sudden bursts of insight that shift reasoning strategies, happened in DeepSeek-R1, which are usually marked by linguistic cues, e.g., “Wait, wait. That’s an aha moment I can flag here.”. The “aha moments” occur when models abruptly recognize the flawed logic, which align with the creative restructuring of human cognition for self-correction. Figure 39 display the number of "aha moments" in DeepSeek-R1 across different NPC problems, where the blue and the red dots represent correct and wrong outputs respectively.
Figure 39: Number of aha moments in DeepSeek-R1
Appendix K Solution Errors
This section visualize the solution errors of different LLMs on the 12 core NPC problems, revealing variations in error distribution across models and difficulty levels. For each problem, each color corresponds to a specific error type as listed in Table 5.
Figure 40: 3SAT
Figure 41: Vertex Cover
Figure 42: Superstring
Figure 43: QDE
Figure 44: 3DM
Figure 45: TSP
Figure 46: Hamiltonian Cycle
Figure 47: Bin Packing
Figure 48: 3-COL
Figure 49: Min Sum Square
Figure 50: Max Leaf Span Tree
Figure 51: Bandwidth
Appendix L Analysis of Reasoning Failure Cases
DeepSeek-R1.
Taking Deepseek-R1 as an example, the reasoning content of its failure cases shows several patterns that lead to wrong answers. The concrete examples of failure cases for DeepSeek-R1 are shown in Table 20. Specifically, the typical reasoning failure cases include:
  • Cascading assumptions and placeholder/guess: Generally, DeepSeek-R1 adopts the trail-and-error approach to find the answer. Although it may give the high-level approach in the beginning, after the first answer is not correct, it will start making assumptions to derive an answer without logical deduction. The following verification may not consider all the given conditions; thus, the verification and the answer are wrong. If it has already made multiple assumptions and verifications without deriving the correct answer, it will return a placeholder or an educated guess.
  • Manual computations errors and complexity avoidance: DeepSeek-R1 makes manual computations that are time-consuming and token-consuming instead of using programming, such as manually counting the length and sum of an array by listing the elements one by one. There are also computation errors that the counting, summation, or sorting results are wrong. In addition, although it already gives the right method, it avoids computation due to complexity and starts making assumptions to simplify the reasoning, which may lead to the incorrect solutions.
  • Reliance on prompt examples: There are examples with the same requirement in the prompt, and DeepSeek-R1 analyzes and generates the answer heavily relying on the example solution, which can be a distraction that DeepSeek-R1 wastes time verifying and editing the example solution.
  • Reasoning repetition: The same assumption, logic, or example can be repeated multiple times without making further progress in solving the problem, which is a waste of time and tokens.
Table 20: Detailed Analysis with Examples of Failure Cases for DeepSeek-R1.
Detailed Failure Reason Reasoning Content
cascading assumptions … Alternatively, perhaps the solution includes nodes 8, 2, 3, 6, 7, 9, 10, 12, 13, 14, 18, 19, 20, 22, 24, 25, 30, 34, 35, 37. Let’s count these. That’s 20 nodes. … Maybe the solution is to select nodes 0, 2, 3, 5, 6, 7, 8, 9, 10, 12, 13, 14, 18, 19, 20, 22, 24, 25, 30, 34, 35, 37. Wait, but that’s more than 20. Need to pick 20. … Wait, perhaps the solution includes nodes 2, 3, 6, 7, 8, 9, 10, 12, 13, 14, 18, 19, 20, 22, 24, 25, 30, 34, 35, 37. Let’s check: …
placeholder/guess … At this point, I’m not sure. Given the time I’ve spent and the complexity, perhaps the solution is the string pyeojmmykmadslpeetdb with length 20, even though some required substrings might be missing. But according to the problem statement, there is a solution. So I’ll have to take a guess. … But given the time I’ve spent and the need to provide an answer, I’ll make an educated guess…
manual computation errors Now, let’s count the number of elements. Each comma separates elements. So between each number there’s a comma. Let’s count commas. But maybe it’s easier to split by each element. Let’s do that step by step. Start from the first element: 1. 1; 2. 14; 3. 1; 4. 1; 5. 5; … So there are 80 items.
complexity avoidance So, if a node is in an edge that’s not covered by any other node, then that node must be in the vertex cover. But checking that for each edge would be time-consuming. Let’s look for edges that have a unique node. For example, if there’s an edge (a, b), and neither a nor b are connected to any other nodes, then one of them must be in the cover. But in a graph of 40 nodes, this is unlikely. So probably most nodes have multiple edges.
reliance on prompt examples … Wait, in the example problem, the solution included nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. So maybe similar patterns here. But the new problem’s edges are different. … Alternatively, perhaps the solution is similar to the example. The example’s solution includes nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. Looking at the problem’s edge set, perhaps some of these nodes are also high-degree here. Let’s see. … Wait, another idea: the example solution included nodes 0, 2, 4, 5, 6, 7, 8, 9, 13, 14, 15, 19, 20, 23, 25, 26, 30, 32, 34, 37. If I look at these nodes, perhaps they are all the nodes from one side of a bipartition. For example, if the graph is bipartite and these nodes form one partition.
reasoning repetition (example) Looking at the edges, let’s see which nodes are involved a lot. For example, node 8 appears in edges like (8,18), (8,35), (8,21), (8,30), (8,39), etc. Maybe node 8 is a hub. … In the problem to solve, node 8 is involved in many edges, as per the edges provided. So including 8 in the cover would cover a lot of edges. … For example, node 8 appears in a lot of edges. Including node 8 would cover many edges. … For example, many edges involve node 8. So including node 8 is a must.
reasoning repetition (logic) … So perhaps the key is to find the longest possible superstring that includes as many of these as possible, then add the ’b’ and the other small strings. But the ’b’ is a problem. Let’s think: the superstring must include ’b’ as a substring. Since none of the other strings contain ’b’, we have to include the string ’b’ in the superstring. So the superstring will have the ’b’ as a standalone. But how? … So in the current problem, perhaps there’s a way to find such a chain where each string overlaps with the next, and ’b’ is inserted in the right place. But I’m stuck. Let’s try to look for possible overlaps between the strings, especially those that include ’mmy’ and ’b’. Wait, the ’b’ must be present. How can ’b’ be included? Since it’s a single character, perhaps it’s placed between two other characters in the superstring. For example, if there’s a string ending with ’a’ and another starting with ’c’, but ’b’ is inserted between them. But in this case, the other strings don’t have ’a’ or ’c’ next to ’b’, so the ’b’ would have to be a standalone. So the superstring must contain ’b’ as a separate character. … So, perhaps the superstring is a combination of several merged strings, with ’b’ inserted in between.
Claude.
Unlike Deepseek-R1, Claude’s failure cases’ reasoning content is generally much shorter. Claude usually starts with giving the high-level step-by-step solution but does not do detailed calculation and verification. It derives a final answer with approximation calculation and claims that the answer is verified as correct. An example is shown in Table 21.
Table 21: Claude Failure Case.
Based on my analysis of the edge patterns, the following nodes would form an effective vertex cover:
- 11 (connects to many nodes)
- 14 (high degree, covers many edges)
- 10 (connects to multiple nodes)
- 16 (appears frequently in edge list)
- 19 (connects to several nodes)
This selection ensures that every edge in the graph has at least one endpoint in our subset, with a total of 11 nodes which meets our constraint.
QwQ-32B.
[…]
Table 22: QwQ-32B Reasoning Representative Examples
Example Reasoning Content
Problem in Format Following … All are present. So this should be a valid solution. Let me confirm again. Another check: For X=5, we chose triple8 (5,13,29). That uses Y=13 and Z=29 … So correct. Yes, this seems to work. So the solution is the list of these triples: {"solution": [[0,10,21],[1,15,28], [2,16,26],[3,14,20],[4,11,23], [5,13,29],[6,18,24],[7,19,25],…]}
Self-Correction in Format Following … Final Answer {"solution": [[3,19,21],[6,13,20],[9,11,26], [4,15,24],[8,10,25],[5,16,23],[7,18,29], [0,17,22],[1,12,28],[2,14,27]]}. Wait, the user’s instruction says to present the answer in the specified JSON format. I think the correct way is to just output the JSON without any extra text. So the final answer is: json{"solution": [[3,19,21],[6,13,20],[9,11,26],…]} Final Answer json {"solution": [[3,19,21],[6,13,20],[9,11,26],…]}
Infinite Reasoning Loop … All are present. So this is a valid solution. The solution is the list of these triples. Final Answer {"solution": [[0,13,29],[1,12,28],[2,17,20], [3,10,26],[4,18,27],[5,14,23], [6,11,25],[7,15,24],…]} Final Answer {"solution": [[0,13,29],[1,12,28],[2,17,20], [3,10,26],[4,18,27],[5,14,23],…]} Final Answer {"solution": [[0,13,29],[1,12,28],…]} … (repeated output continues)
Appendix M Costs of the Evaluation
Table 23 displays the number of input token and completion token with their corresponding prices, and the total cost of running online models once for all difficulty levels across core NPC problems.
Table 23: Cost for online models
Model Prompt Completion Cost
GPT-4o-mini 30964144 (\$0.15/MTok) 9442548 (\$0.6/MTok) \$10.31
GPT-4o 30963606 (\$2.5/MTok) 7786156 (\$10/MTok) \$155.27
Claude-3.7-Sonnet 33799101 (\$3/MTok) 11186272 (\$15/MTok) \$269.19
DeepSeek-V3 31490957 (2RMB/MTok) 16178388 (8RMB/MTok) 192.41RMB
DeepSeek-V3-2503 31490957 (2RMB/MTok) 31808451 (8RMB/MTok) 317.45RMB
DeepSeek-R1 31512557 (4RMB/MTok) 95936418 (16RMB/MTok) 1661.03RMB
o1-mini 31360984 (\$1.1/MTok) 35161551 (\$4.4/MTok) \$189.21
o3-mini 31199884 (\$1.1/MTok) 110944621 (\$4.4/MTok) \$522.48